A Complete Bibliography of Publications in the journal Mathematics in Computer Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

28 July 2025 Version 1.34

Title word cross-reference

(a, b, c) [230]. 1 [288]. 10 [449]. 2 [88, 198, 508]. 3 [320, 372, 106, 155]. 3n [384]. 3 × 3 × 3 [251]. 8p [618]. C_{6} [590, 592]. D [133, 583]. H^{1} [515]. K [233, 110, 111, 547, 78, 80, 217]. K[X][Y; f \frac{d}{dX}] [480]. K_{m,n}[P_{k}] [390]. L(2, 1) [146, 283]. m [341, 180]. S_{6} [419]. n [181, 504]. Ω [52]. π [62]. q [579]. SL_{3}(C) × SL_{3}(C) × SL_{3}(C) [251]. T [329]. τ [86, 336]. ε [216]. W [374].

-Adic [86]. -Algebras [233]. -Analogs [579]. -approach [216]. -arboricity [547]. -Body [384]. -Calculus [62]. -Coloring [329]. -Fault [288]. -Fixed-Endpoint [80]. -Fold [180]. -Labeling [146, 283]. -Means [217]. -Method [336]. -Module [133]. -Modules [583]. -Orthogonal [508].

-Phase [78]. -Polar [341]. -SAT [110, 106]. -sectors [504]. -Trees [111, 155]. -Year [449].

2-Orthogonal [604]. 2D [425].

3 [191].

4.0 [549].

84 [622].

Aachen [570]. Abelian [488, 547]. Absent [537]. Absent-Minded [537]. Abstraction [578]. ACC [197]. ACC/DEC [197]. Accelerating [240]. Accelerations [198]. Access [586]. Accuracy [333]. Accurate [434, 173, 487]. Achievements [435]. Activity [543]. Adams [465]. Adic [86]. Adjoint [257]. Advanced [442, 270, 362].

Advances [335, 392, 506]. AES [89]. Affine [401, 462, 257]. After [347]. AGE [281]. Agents [553]. Agreeing [231]. Ahead [197]. Aircraft [446]. Algebra [576, 572, 15, 560, 596, 432, 486, 564, 434, 65, 479, 476, 241, 449, 578, 383, 537, 542, 622, 583]. Algebraic [70, 593, 139, 519, 118, 548, 257, 483, 481, 350, 160, 454, 368, 117, 164, 140, 28, 190, 452, 13, 223, 63, 256]. Algebraically [302]. Algebras [233, 367, 482, 339, 395, 509, 379, 308, 232]. Algebro [551]. Algebro-Geometric [551]. Algorithm [236, 455, 401, 123, 529, 575, 219, 257, 373, 106, 531, 469, 69, 357, 388, 105, 80, 441, 263, 231, 527, 281]. Algorithmic [613, 307, 534, 32]. Algorithms [593, 37, 146, 302, 350, 107, 71, 193, 524, 103, 63, 59]. Alignment [219]. Alignment-Based [219]. Alignments [539, 400, 362]. Allowing [30]. Almost [287, 619, 550]. almost-circles [550]. Alternative [19]. Ambient [305]. Amoebas [597]. Amounts [312]. Analogs [579]. Analysing [476]. Analysis [110, 393, 425, 279, 477, 212, 36, 303, 167, 200, 246, 268, 77, 492, 164, 338, 440, 28, 541, 556, 380, 210, 271, 551, 217]. Analytic [12, 126, 7]. Analytical [108, 502]. Analyze [485]. AND/OR [59]. Andreas [517]. Angles [504]. Animated [375]. Animations [435]. Anisotropically [377]. Annotated [172]. Anonymization [405]. Antimagic [226, 286, 390, 151, 225]. Antimagicness [282]. Antiprisms [226]. AntiSARS [553]. AntiSARS-CoV2 [553]. Ants [365]. Any [471]. Applicability [424]. Application [341, 465, 212, 535, 25, 303, 606, 214, 447, 181, 263, 270, 383]. Applications [285, 536, 116, 144, 575, 371, 8, 74, 36, 595, 609, 325, 32]. Applied [340, 511]. Approach [127, 12, 227, 167, 31, 168, 21, 447, 307, 128, 602, 396, 65, 252, 196, 382, 452, 138, 526, 542, 218, 137, 216]. Approaches [28]. Approximate [158, 227, 21, 456, 6, 137]. Approximating [349]. Approximation [5, 22]. Approximations [74, 602]. Arbitrarily [148]. Arbitrary [125, 504]. arboricity [547]. Archimedean [202]. Area [155]. Arithmetic [171, 190, 476, 260]. Arrangements [115, 116]. Array [357]. Arrays [415, 358, 546, 545]. Art [144, 385]. Artifact [564]. Artificial [553]. Aspects [586, 597]. Assignment [37]. Assisted [542]. Associated [410, 383, 527, 583]. Associativity [342]. Astronomy [424]. Asymptotically [372]. ÁTAME [495]. ATP [277, 51]. ATP-based [51]. Attacks [407]. Attractive [447]. Attractor [59]. Atwood [380]. Authoring [92]. Automata [34, 176, 204, 364]. Automated [61, 244, 500, 429, 580, 503, 440, 501, 505]. Automatic [471, 50, 435, 317]. Automatically [203]. Automaticity [348]. Automating [344]. Automorphism [489, 419, 590, 592]. Autonomous [519, 481]. Average [77, 359]. Averaged [479]. Avoid [255]. Avoiding [407, 601]. Aware [207]. Axis [196].

B [222]. B-Spline [222]. Backward [353]. Bad [93]. Ball [195]. Ball-end [195]. Banach [6]. Banana [148]. Band [514]. Bar [575, 485]. Bargraphs [544]. Base [413, 221]. Based [285, 219, 485, 97, 302, 469, 313, 266, 408, 524, 140, 378, 85, 196, 271, 526, 557, 362, 197, 217, 327, 137, 213, 214, 281, 51, 369]. Bases [238, 19, 460, 87, 623, 158, 566, 467, 521, 531, 532, 522, 619, 610, 68, 422, 346, 10]. Bashforth [465]. Basic [437]. Basics [405]. Basis [469, 69, 375, 186, 441]. Battle [588]. Baxter [131, 132]. Bayesian [67]. be [310, 232]. BE-algebras [232]. Behavior [70, 221, 60]. Behaviour [86]. Berge [295]. Bernstein [263]. Bertini [611]. Bertrand [616]. Beta [558]. Between [391, 183, 346, 507]. Beyond [549]. Bezout [22]. Bi [451]. Bi-univalent [451]. BIBD [414]. Bibliographic [585]. Bifurcation [63]. Bifurcations [13]. Big [405]. Bill [39]. Binary [425, 489, 220, 360, 327]. Binomial [121, 448]. Biochemical [61]. Biological [534, 164, 165, 28]. Biology [63]. Bioregulatory [60]. Bipartite [283, 327]. Birkhoff [422]. Bisection [397]. Bispanning [73]. Bit [105]. Bit-Parallel [105]. Bitstream [138]. Bivariate [161, 437, 610]. Bivium [87]. Black [563]. BLIM [105]. Block [618, 361]. Blocking [108]. Blocks [413]. BLS [558]. Bodies [530, 560, 596]. Body [377, 474, 384]. Boltzmann [587]. Book [381]. Boolean [411, 623, 27, 60, 59]. Border [467, 469]. Bound [262]. Boundary [450, 223]. Bounded [450, 198]. Bounding [455]. Bounds [521, 252, 267, 451]. Bow [292]. Box [269]. Boxes [563, 275, 262]. Branch [374, 262]. Branch-and-Bound [262]. Breaking [457]. Breaks [514]. Brief [382, 404]. Bring [578]. Brylawski [185]. Buchberger [373, 346]. Burgers’ [281, 121]. Burrows [359]. Butson [421].

C [172, 563, 169]. Cache [107]. CAD [117, 608]. Calculating [593]. Calculation [497, 14, 376, 475, 448]. Calculator [476, 578]. Calculus [207, 62, 175, 127, 440, 442, 223, 222]. CAM [193, 196]. Can [578, 310]. Canal [258]. Canny [599]. Canonical [323, 12]. Capable [273]. Caputo [512]. Carleman [340]. Carlitz [554]. Cartesian [151, 227, 41]. CAS [462, 371, 430, 485, 378]. CAS/DGS [485]. Case [379, 248, 77, 441, 442, 359, 608]. Cases [594]. Catalan [601, 544]. Category [367, 482]. Cauchy [127]. Caylay [547]. Cell [438, 273]. Cells [555]. Centers [318]. Central [372, 384]. Centroid [285]. Certain [562, 295, 293, 300, 451, 329]. Certified [318]. Certify [15]. Chains [123, 518, 331]. Challenge [537]. Challenges [434, 435, 405]. Challenging [319]. Change [377, 344, 587]. Channel [407, 444]. Chaotic [86]. Characteristic [409, 88, 91, 346]. Characterization [264, 418]. Chebyshev [333, 507]. Checking [317, 304]. Chemical [166, 228]. Choice [244]. circles [550]. Circulants [300]. Circular [530, 596, 382]. Class [261, 384]. Classes [334, 511, 461, 316, 240]. Classical [391, 250]. Classification [480]. Classifying [350]. Classroom [576]. Classrooms [375]. Clause [304]. Claw [145]. Claw-Free [145]. Clipping [161]. Clique [154]. Clocks [46]. Closed [302, 150, 331]. Closed/Open [150]. Cluster [36]. Clustering [523, 217]. Clusters [237, 229, 21]. CNC [193]. Coarseness [143]. Coboundary [183]. Cocyclic [618, 416]. Code [593, 575, 352]. Codes [417, 488, 410, 489, 198, 156, 189, 186, 85, 461, 327]. Coding [157]. Coefficients [121, 451, 398, 10, 448]. Cohen [379]. Collaborative [247]. Collection [96]. Collections [98]. Colliers [458]. Coloring [78, 329, 459]. Column [554]. Column-Convex [554]. Combinations [301]. Combinatorial [418, 546, 545]. Combining [200]. Common [30, 365, 364, 538, 507]. Communication [104]. Communications [586]. Community [24]. Commuting [619]. Companion [428]. Comparison [491, 188, 32]. Complete [575, 169]. Completely [11]. Completeness [216]. Complex [450, 420, 419, 523, 17, 513]. Complexity [104, 228, 66]. Component [207]. Components [559]. Composite [516, 422]. Composition [207]. Compositional [539]. Compositions [555, 463]. Comprehension [496]. Comprehensive [302, 439, 583]. Compressed [358]. Compression [104, 220, 219]. Computation [427, 124, 334, 340, 511, 379, 531, 335, 392, 506, 280, 394, 604, 386, 67, 4, 608].

Computational [36, 135, 66, 597, 338, 516, 446]. Computationally [485]. Computations [566, 530, 375, 140, 40]. Compute [494, 257, 602]. Computer [576, 572, 15, 560, 596, 432, 486, 564, 434, 347, 105, 479, 241, 449, 578, 383, 537, 622, 276]. Computing [407, 552, 192, 129, 323, 229, 114, 158, 467, 562, 619, 69, 330, 312, 186, 474, 91, 439, 181, 441, 103, 174, 527, 583, 611, 3, 10]. Concept [38]. Concepts [385]. Concretizations [116]. Conditional [152, 178, 81]. Conditions [145, 8, 606, 516, 589]. Cone [382]. Cones [109]. Conference [421]. Configuration [485]. Configurations [588, 383, 384]. Conformal [320, 135]. Conforming [515]. Confront [442]. Conic [5, 382]. Conjecture [615, 295]. Conjecturing [576]. Conjugation [340]. Connected [530, 560, 596, 299, 82]. Connection [346]. Connectivity [328, 309]. Consensus [615]. Conserved [613]. Constants [568, 4]. Constrained [459]. Constraint [272]. Constraints [301, 261]. Constructibility [317, 316]. Constructible [323]. Construction [285, 579, 320, 491, 419, 357, 77, 479, 151, 461, 317, 359]. Constructions [488, 327]. Constructive [133]. Containing [590, 592]. Context [100, 97]. Context-Free [97]. Continuation [255, 7]. Contractions [339]. Contribution [518, 404]. Control [234, 393, 584, 447, 313, 224, 337, 273, 197]. Controlled [444]. Convergence [15, 161]. Convergent [139, 439]. Conversion [351, 472]. Convex [192, 450, 303, 554, 305]. Coq [279, 491, 248]. Coquelicot [279]. Core [615]. Corona [282]. Coronene [290]. Corpus [317]. Correcting [156]. Correction [546, 590]. Cost [77]. Counterpart [581]. Counting [310, 588, 544]. Coupled [467]. CoV2 [553]. Cover [426, 75, 80, 154]. Coverage [178].

Covering [401, 212, 227, 546, 545, 178, 81]. Coxeter [468]. Cracked [338]. Creative [128, 533]. Credibility [605]. Cross [51]. Cryptanalysis [90, 85]. Cryptographical [188]. Cryptography [408]. Cryptosystem [90]. Cryptosystems [85]. CSP [266]. CTL [393]. Cube [82]. Cube-Connected [82]. Cuckoo [77]. Curvature [197]. Curvature-Based [197]. Curve [113, 88, 257]. Curved [192]. Curves [407, 70, 118, 462, 548, 600, 486, 368, 590, 592, 558, 91, 408, 607, 306, 324, 137, 222]. Cutter [195]. Cycle [188, 293, 187, 299]. Cycles [82]. Cyclic [488, 85, 589]. Cylindrical [160, 454, 256].

D [320, 372, 198]. DAEs [445]. Dags [140]. Data [20, 536, 424, 235, 36, 47, 405]. Database [493]. Databases [404]. Dataflow [44, 41]. Datasets [608]. Date [471]. DC [452]. Dealing [398]. DEC [197]. Decides [205]. Decision [204]. Declarative [44]. Decomposition [349, 160, 454, 396, 186, 508]. Decompositions [438, 538, 256]. Dedication [517, 528]. Deduction [500, 501, 505]. Defined [409, 137]. Definition [376, 475]. Degenerate [363, 31]. Degree [145, 411, 287, 521, 547, 558, 187]. Degrees [43]. Delay [465]. Delayed [393]. Delegation [498]. Dendrimer [328]. Density [38]. Dependence [260]. Dependent [515, 272]. Deprit [479]. Depth [463]. DERIVE [449]. Deriving [60]. Description [213, 55]. Description-based [213]. Descriptive [391]. Design [385, 168, 564, 388, 140, 446]. Designing [553]. Designs [414, 579]. Detect [370]. Detecting [324]. Detection [613, 503, 59]. Determination [265, 267]. Deterministic [611]. Devices [434]. DG [196]. DG/K [196]. DG/K-Based [196]. DGS [462, 485]. Diagrammatic [471]. Diagrams [552, 464]. Diamond [294].

Dichotomy [27]. Dickey [619]. Difference [415, 413, 483, 613, 437, 484, 565, 610]. Differences [548]. Different [188, 377, 474, 476, 461]. Differential [465, 123, 477, 124, 481, 379, 535, 333, 132, 522, 201, 122, 268, 120, 437, 484, 603, 399, 510, 604, 607, 126, 617, 526, 11, 125, 527, 336, 10]. Differentiated [616]. Differentiation [181]. Diffusion [535, 353]. Digital [425, 93, 442]. Digitalization [92]. Digitization [94, 95]. Digraphs [287]. Dimension [488, 519, 483, 120, 437, 484, 565, 610, 179, 294, 252, 125]. Dimensional [368, 602, 338]. Dimensionality [525]. Directed [296]. Directions [144]. Directly [376, 475]. Disambiguation [56]. Discovering [203]. Discovery [36, 215]. Discrete [202, 493, 371, 164, 198, 121, 344, 587, 222]. DiscreteZOO [493]. Discretization [281]. Discriminant [253]. Disease [64]. Diskew [367]. Distance [593, 433]. Distances [285]. Distributed [269, 31]. Distribution [601]. Distributions [237]. Divisions [575]. Divisive [426]. Divisors [69]. DNA [236, 355, 31]. DNA/RNA [31]. Do [370, 568]. Documents [49, 100]. Does [370]. Domain [245]. Domains [581]. Dominating [330]. Domination [298, 290]. Drawings [155]. Driven [247]. Dual [486, 389]. Duality [185]. due [514]. Duopoly [616]. During [442]. Dynamic [427, 315, 432, 486, 447, 375, 435, 381, 382, 585, 502]. Dynamical [167, 27, 203]. Dynamics [62, 596, 337, 60].

Earth [530]. Ebola [337]. Eccentric [328]. Economics [447]. Edge [242, 289, 149, 299]. Edge-Magic [149]. Editions [95]. Editorial [1]. Education [549]. Effect [476]. Effective [7]. Effectively [520]. Effects [516]. Efficiency [330, 108]. Efficient [236, 495, 455, 37, 45, 498, 169, 303, 261, 559, 388, 592, 386, 190, 590]. Efficiently [520]. Eigenfunctions [604].

Eighth [162]. Einstein [483]. Electrification [452]. Electro [434]. Electro-optic [434]. Electrostatic [383]. Element [515, 472]. Elements [577, 562, 434]. Elimination [166, 21, 126, 304, 26]. Ellipsoid [447]. Elliptic [407, 408]. Embedding [558, 165, 291, 300]. Emiris [599]. Empirical [20]. Employment [514]. Encapsulating [245]. Encyclopedia [318]. end [195]. Endpoint [80]. Energy [108]. Energy-Time [108]. Enforcement [495]. Engineering [357, 247, 505]. Enhance [330]. Entangled [422]. Enthusiast [572]. Enumerating [557]. Enumeration [131, 132]. Envelope [435]. Envelopes [548, 580]. Environment [427]. Enzyme [362]. Epidemiological [64]. Epidemiology [65]. Equal [348, 79]. Equal-Length [79]. Equalising [584]. Equation [515, 594, 114, 604, 121, 281]. Equations [465, 129, 139, 477, 89, 124, 483, 481, 535, 25, 606, 180, 399, 607, 617, 526]. Equi [422]. Equi-Entangled [422]. Equilibria [64]. Equilibrium [530]. Equivalence [238, 174]. Error [156]. Error-Correcting [156]. Errors [35, 56, 442]. Estimating [109, 514]. Estimation [200, 274]. Euclid [428]. Eulogy [403]. European [551]. Evader [268]. Evaluating [383]. Evolution [137, 96]. Evolution-Based [137]. EX [83]. Exact [25]. Exam [578]. Examples [376, 475]. Exchange [187]. Exercises [101]. Exist [568]. Existence [27, 421]. Exists [261]. Expanded [54]. Expanding [108]. Expansions [86, 451]. Experiments [546, 545, 51]. Explain [543]. Explicit [445, 510]. Exploration [548, 580, 486]. Exploring [334, 462]. Exponential [445, 436]. Exponentiation [498, 621]. Expressed [19]. Expression [140]. Extend [57]. Extended [158, 344]. Extension [465, 480, 420, 190, 293]. Extensions [120, 437]. External [357, 359]. Extrapolation [66]. Extreme [605].

Faber [451]. Face [226, 225]. Factor [125]. Factorable [11]. Factorials [546, 545]. Factoring [563]. Factorization [33, 71, 122]. Factorizations [363, 11, 125]. Factors [538]. Familiarizing [376, 475]. Families [230, 413, 188, 297, 187, 551]. Family [604, 583]. Fan [109]. Fans [234]. Fast [128, 375, 103]. Faster [558, 359]. Fault [418, 288]. Feller [353]. FEM [595]. Fermat [462]. FEVS [174]. Fiber [221]. Field [302, 560, 120, 437, 308, 383]. Fields [409, 120, 190]. Filtered [373]. Filters [232]. Finding [74, 526]. Fingerprint [493]. Finite [409, 577, 613, 434, 260]. Finite-Difference [613]. Finitely [332, 120]. Fire [516]. First [481, 205, 322, 617, 51]. First-Order [205]. Five [287]. Fixed [148, 27, 80, 557]. Fixed-Point [27]. Flatness [200]. Flow [444, 210]. Flows [606]. Focus [235, 500, 301, 193, 335, 392, 506, 325, 276]. Fold [180]. Folded [291, 153]. Following [139]. Forall [261]. Foreword [58, 163, 92, 423, 470, 29, 235, 500, 102, 366, 453, 249, 517, 528, 561, 591, 170, 84, 301, 443, 112, 412, 354, 48, 243, 23, 182, 490, 193, 335, 392, 506, 402, 72, 141, 284, 208, 325, 259, 199, 119, 573, 314, 276, 134, 9, 2]. Form [123, 478, 562, 195]. Formal [612, 172, 245, 246, 307, 276]. Formalization [239, 378]. Formalized [491]. Format [53]. Forms [614, 431, 527]. Formula [599, 97, 181]. Formulae [510]. Formulas [184]. Foundations [218]. Four [287, 590, 592]. Fourier [307]. Fourth [11]. Fractal [425]. Fractals [549, 189, 543]. Fraction [538]. Fraction-Free [538]. Fractional [512, 535, 353]. Fragment [205]. Framework [115, 418, 369, 605]. Francs [159]. Free [145, 111, 211, 64, 97, 373, 213, 195, 538].

Free-form [195]. Freedom [57].

Frequencies [157]. Friendly [279, 558]. Fuel [273]. Fulkerson [295]. Full [417]. Fullerene [242]. Function [264, 175, 395, 509, 623, 268, 181, 440]. Functional [130, 45, 398, 174]. Functions [450, 272, 581, 411, 12, 131, 132, 602, 203, 91, 126, 451, 3]. Fundamental [251]. Fundamentals [135]. Future [144, 586]. Fuzzy [234, 233, 341].

G01 [198]. Gain [167]. Galois [472]. Game [616]. Games [268]. Gaussian [237, 21]. GCD [456]. Gelfand [619]. General [512, 115, 498, 38, 598, 606, 201, 396, 604, 138]. Generality [370]. Generalized [367, 282, 185, 289, 148]. Generated [120]. Generating [416, 131, 132, 606]. Generation [471, 195]. Generic [207, 140]. Genetic [350]. Genus [88, 590, 592]. Geodesic [321]. GeoGebra [585]. Geometer [540]. Geometric [576, 239, 303, 368, 184, 241, 317, 551]. Geometries [211]. Geometry [536, 427, 315, 500, 135, 432, 486, 375, 435, 568, 381, 501, 241, 382, 585, 319, 446, 502, 543]. Gerdt [574, 572, 528, 569, 571, 570, 567]. Get [56]. GGH [90]. Given [547]. Global [512, 15, 21, 263, 217]. Globally [139]. Gluing [231]. Goethals [413]. Good [70, 93, 468]. Graceful [144, 148, 292]. Grammar [97]. Grammars [308]. Grammatical [97]. Granulations [217]. Graph [285, 584, 50, 372, 229, 457, 78, 390, 308, 328, 278, 325, 296, 225, 154, 459]. Graph-Based [285]. Graph-Modelled [584]. Graphical [487]. Graphs [341, 286, 73, 326, 143, 145, 282, 38, 75, 361, 621, 71, 227, 76, 147, 289, 292, 547, 433, 188, 80, 311, 283, 151, 293, 298, 300, 288, 210, 297, 178, 187, 299, 83, 327]. Graver [186]. Gravitational [560]. Greedy [575, 599]. Grid [296, 557]. Grid-Based [557]. Gröbner [460, 109, 87, 623, 158, 302, 351, 532, 522, 69, 375, 610, 68, 439, 583, 346, 10]. Groebnerialium [574]. Grothendieck [441]. Ground [205]. Group [88, 590, 592]. Groups [498, 547]. Guard [266]. GVW [531].

Hadamard [415, 618, 416, 417, 420, 419, 422, 421]. Half [552, 445]. Half-Explicit [445]. Half-Lines [552]. Hamiltonian [612, 288]. Hammer [277]. Harmonic [581]. Hashing [77]. Having [489]. Heat [541]. Heater [234]. Height [428]. Helicopter [445]. Help [430, 365]. Hereditary [142]. Hermite [595, 436, 473]. Hermitian [431]. Heuristic [416]. Hexagonal [152]. Hierarchical [99]. Hierarchies [619]. High [255, 577, 535, 174, 487]. High-Order [577, 535]. High-Performance [174]. Highly [355, 553]. Historical [47]. Hit [55]. HIV [393]. HOL [277]. Holonomic [583]. Homogeneous [253]. Homological [252]. Homology [464]. Homotopy [255]. Honeycomb [152]. Honour [39]. Hopf [63]. Hori [479]. Hough [536, 324]. Huffman [157]. Hull [192]. Hybrid [62, 176, 303, 266, 247]. Hydropathy [362]. Hyperbolic [25, 587]. Hyperbolicity [466]. Hypercubes [291, 153]. Hyperdeterminant [251]. Hyperelliptic [590, 592]. Hypergeometric [180]. Hypergraph [342, 177]. Hypergraphs [142]. Hypersurface [527, 583]. Hypersurfaces [305].

ICT [369]. Ideal [410, 257, 559, 69, 439, 386]. Ideals [532, 562, 179, 611]. Identifiability [265]. Identification [492]. Identify [236]. Identifying [49]. II [116, 86, 392]. III [506]. IL-System [26]. Image [220, 212, 214, 492]. Images [425, 324]. Immune [393]. Impact [449, 585]. Implementation [45, 563, 374, 388, 140, 479, 331].

Implementations [220]. Implicit [70, 477, 114, 181, 526]. Implicitly [137]. Improve [454]. Improved [146, 219, 77, 263, 231]. Improving [333]. Inaction [514]. Incidence [512]. Indefinite [240]. Indeterminate [363]. Index [445, 328]. Index-1 [445]. Indicator [188, 187]. Individuals [65]. Indo [551]. Indo-European [551]. Industry [263]. Inequalities [576]. Inference [269, 26]. Infinitary [43]. Infinite [40]. Inflection [442]. Inflexions [600]. Influence [424]. Information [285, 584]. Infrastructure [115]. Initial [451]. Inner [267]. Input [444]. Insertion [308]. Insights [553]. Instances [110]. Instead [3]. Instructor [542]. Instructor-Assisted [542]. Integrable [594]. Integral [334, 340, 511, 399, 439, 376]. Integrals [369, 617]. Integrating [44]. Integration [465, 46]. Integro [510]. Intelligence [553]. Intelligent [447, 381]. Intensional [45, 41]. Interactive [100, 101]. Interesting [503]. Intermediate [120]. Interplay [391]. Interpolation [595, 564, 436, 473, 198]. Interpolator [197]. Interpretation [268]. Intersecting [414]. Intersection [266]. Interval [264, 200, 268, 80, 178, 273, 270, 271, 17]. Intervals [275]. Intracellular [393]. Introducing [584, 509]. Introduction [150, 209]. Intuitionistic [234, 233]. Invariant [395, 509, 368, 447, 184]. Invariants [251, 611]. Inversion [12]. Invertible [524]. Investigating [63]. Investigation [612, 596, 368, 622]. Involutive [575, 521, 531, 532]. Involving [272]. Irregularity [242, 289, 288, 297]. Isabelle [278]. Isodual [489]. Isogenous [407]. Isolated [527, 583]. Isolating [138]. Isolation [16]. Isomorphic [407]. Isoptics [462]. Isoscallop [195]. isotropically [474]. Isotypeness [238]. Issue [92]. Items [350]. Iteration [41]. Iterative [139, 162].

Jahangir [225]. Janet [574]. Jerk [224]. Jobs [79]. Join [286]. Journal [620]. Jugs [312].

K-Based [196]. Kernel [264, 340, 511, 541, 293]. Kernels [564]. Key [89, 187]. Kids [543]. Kind [507]. Kinds [476]. Kinetic [25]. Kites [589]. Kleisli [491]. Knowledge [238, 52, 92, 235, 239, 343, 378, 46, 501]. Krull [120]. Krull-type [120]. Kummer [88]. Kutta [445].

Labeling [341, 146, 390, 283, 151, 225]. Labelings [230, 226, 286, 76]. Labelling [144]. Labor [514]. Lambert [374]. Lanczos [336]. Landscapes [110]. Lane [491]. Language [54, 551]. Languages [45, 245]. Large [188, 187, 98]. Larger [91]. Last [601]. Latin [485]. Lattice [544, 294, 587]. Lattices [184]. Law [88]. Layout [296]. Lazy [45]. Learn [430]. Learning [371, 454, 442, 608]. Lebesgue [376, 475]. Legs [310]. Lempel [33]. Length [348, 79]. Less [33, 348]. Lessons [608]. Level [268, 578]. Lexicographic [390]. Libraries [93]. Library [279, 278, 446, 96]. Lie [305, 339, 308]. Liénard [594]. Liénard-type [594]. Light [136, 277]. Lightweight [357]. Likert [350]. Limitation [105]. line [355]. Linear [129, 272, 410, 130, 529, 124, 176, 594, 180, 547, 524, 126, 153, 296, 260, 11, 125, 542, 483]. Lines [552, 47]. Linkage [485]. Liquid [587]. Local [110, 70, 124, 71, 227, 322, 441, 215, 400, 32]. Localization [559, 271]. Locally [145]. Location [316, 17]. Locus [427, 435]. Logic [234, 205, 46, 526]. Logical [63]. Longest [364]. Look [197]. Look-Ahead [197]. Loose [75]. Lorenz [622]. Lorenz-84 [622]. Loss [370]. Lossless [104]. Lotka [368]. Low [104, 564, 602, 22, 359]. Low-Artifact [564]. Low-Complexity [104]. Low-Rank [602]. LTL [304]. Lucid [44, 41]. Lumping [394]. Lyapunov [313, 203]. Lyapunov-Based [313].

MAAOR [394]. Mac [491]. Machine [454, 380, 196, 608]. Machining [198, 196]. Magic [230, 149, 154]. Malicious [498]. Management [92, 501]. Mandala [121]. Manufactured [606]. Manufacturing [168]. Many [568]. Maple [471, 563, 430]. Mappings [524]. Maps [471, 187, 13]. Market [514]. Markov [372, 598, 444]. Markovian [331]. Markup [53]. Masses [377, 474]. Matching [37, 355, 31, 105, 311]. Matching-Star [311]. Math [54, 55]. Mathematica [380, 376, 475]. Mathematical [49, 52, 92, 100, 97, 343, 95, 319, 553, 96]. Mathematics [549, 93, 235, 371, 347, 193, 344, 276]. MathML [191]. Matrices [415, 618, 416, 420, 428, 179, 510, 609, 422, 187, 421]. Matrix [19, 12, 419, 602, 538, 22]. Matroids [184]. Maude [495]. Max [224]. Maximin [150]. McEliece [85]. MDS [461]. Mean [370, 383]. Mean-Value [383]. Meandering [306]. Meandric [458]. Means [564, 479, 451, 217]. Measurable [312]. Measure [213, 475]. Mechanical [172, 540]. Mechanics [505]. Mechanisms [485]. mega [52]. Membership [494, 439, 265]. Membrane [330]. Memories [567]. Memory [357, 359, 570, 459]. Mesh [281]. Messages [269]. Method [515, 401, 129, 424, 61, 529, 333, 434, 266, 399, 479, 587, 611, 502, 336, 398, 51, 353, 16]. Methodology [606, 270]. Methods [465, 478, 535, 560, 596, 66, 245, 162, 445, 164, 394, 273, 241, 270, 622]. Metric [213, 294]. Metrics [211]. Microarray [36]. Middle [74]. Miller [403, 404]. Min [224]. Minded [537]. Minima [74]. Minimal [428, 311]. Minimax [150]. Minimization [623]. Minimum [593, 365, 330, 155].

Minkowski [349]. MINLP [263]. Mirka [403, 404]. Mirroring [229]. Mismatch [35]. Mixed [416, 599, 397]. Mixture [237]. Mizar [51]. Möbius [183]. Mode [274]. Model [393, 512, 166, 237, 25, 67, 270, 513, 331]. Modeling [372, 434]. Modelled [584]. Modelling [553]. Models [238, 61, 64, 598, 164, 149]. Modified [436, 473]. Modular [460, 186, 386]. Module [133]. Modules [480, 373, 583]. Moduli [250]. Molecules [321]. Monochromatic [220]. Monodromy [255]. Monomial [129]. Monomial-by-Monomial [129]. Moon [465]. Moore [287]. Motifs [236]. Motion [224, 479, 380]. Motions [320, 560]. Motivating [537]. Moulton [465]. Moving [530, 560]. Multi [168, 444, 196, 400, 217]. Multi-Axis [196]. Multi-channel [444]. Multi-Granulations [217]. Multi-Objective [168]. Multi-way [400]. Multidimensional [349, 553]. Multilingual [219]. Multinomial [448]. Multiple [586, 533, 534, 514, 47, 331]. Multiplication [179, 280]. Multiplying [488]. Multiprecision [171]. Multitarget [543]. Multivariate [563, 8, 431, 18, 484]. Mutations [30]. My [578].

Nanning [570]. Nanoscale [434]. Nanotechnology [434]. navigation [620]. NC [196]. Near [332, 214, 215, 209, 210, 421, 218]. Nearly [421]. Nearness [212, 213, 214, 216, 217]. Necklace [300]. Neighborhood [150, 217]. Neighbourhoods [213]. Network [406, 290, 331]. Networking [580]. Networks [584, 146, 24, 269, 228, 534, 444, 165, 152, 60, 513, 329, 59]. Newton [529, 494, 113, 114, 254]. Noise [424]. Non [202, 455, 272, 30, 269, 340, 586, 342, 86, 157, 474, 583, 421]. Non-Archimedean [202]. Non-Associativity [342]. Non-existence [421]. Non-isolated [583]. Non-isotropically [474]. Non-Linear [272]. Non-Optimality [86]. Non-overlapping [30]. Non-preemptive [455]. Non-Sorted [157]. Noncommutative [460, 189, 223]. Nonexistence [287]. Nonholonomic [313]. Nonisomorphism [50]. Nonlinear [124, 478, 333, 167, 262, 266, 524, 187]. Nonlocal [127]. Nonplanar [59]. Nontrivial [383]. Nonuniform [178]. Nori [464]. Normal [614, 123, 478]. Note [145, 414, 473, 382]. Nugis [574]. Number [143, 237, 430, 140, 293, 557]. Numbers [439, 3]. Numeric [529, 168, 356, 335, 392, 506]. Numerical [465, 172, 477, 494, 15, 487]. Numerically [3]. NURBS [197].

Object [44]. Object-Orientation [44].

Objective [168]. Objects

[192, 493, 239, 492]. Oblivious [107].

Obtaining [89]. Odd [489, 558]. ODE [61]. ODEs [529, 519]. ODOs [619]. Off [281]. Off-Step [281]. Offsets [70, 548, 580]. Olympiad [319]. On-line [355]. One [143, 519, 348, 395, 268]. One-Parameter [395]. ones [550]. Online [277]. Only [89]. Open [150, 446]. OpenMath [57, 191]. Operational [127, 510, 223]. Operations [303, 77, 386]. Operator [340, 125]. Operators [334, 511, 562, 122, 437, 11, 10]. optic [434]. Optical [246]. Optimal [393, 457, 162, 224, 80, 337, 305]. Optimality [86]. Optimising [87]. Optimization [15, 168, 263]. Optimizing [160]. Orbit [465, 530, 596]. Order [618, 450, 577, 410, 489, 481, 535, 499, 162, 356, 388, 205, 322, 604, 181, 187, 125, 582, 11]. Ordinary [477, 124, 481, 535, 526]. Ore [480]. Organization [52, 513]. Orientation [44]. Orientations [530]. Oriented [293]. Orthogonal [508, 604, 609, 586]. Oscillations [478]. Oscillator [78].

Oshima [487]. Other [147, 589]. Outer [419, 267, 475]. Output [487]. Overcoming [105]. Overlap [391]. overlapping [30]. Overthrow [241]. Oxide [273].

P [569]. Paced [542]. Pagenumber [82]. PageRank [394]. Paired [511]. Pairing [558, 408]. Pairing-Based [408]. Pairwise [424]. Palindromic [364]. Paradigm [44, 362]. Paradigms [608]. Parallel [552, 104, 105, 220]. Parallelepipeds [595]. Parameter [395, 509, 388, 274]. Parameter-Order [388]. Parameterization [258, 436, 137]. Parameters [290, 352]. Parametric [115, 116, 269, 176, 369, 206, 439, 11, 446, 611, 137]. Parametrizations [221]. Parametrized [322]. Part [594]. Partial [124, 485, 227, 122, 603, 11, 125]. Partially [400]. Particles [269]. Particular [160]. Partition [365, 299]. Partitional [76, 147]. Partitioning [73, 229, 306]. Passengers [537]. Path [139, 499, 496, 195, 80]. Path-Following [139]. Pattern [601, 220, 36, 31, 105, 215]. Patterns [248, 438]. PBD [414]. PBW [583]. PDE [524]. Pendulum [478]. Perceptual [214]. Perfect [415, 421]. Performance [174]. Periodic [13]. Permutation [438, 283]. Persistent [464, 387]. Personal [569]. Perspective [433, 449]. Perturbation [25, 322]. Perturbations [474]. Perturbed [507]. Phase [78, 156, 271, 587, 606]. Phylogenetics [551]. Physics [433]. Picard [529]. Piecewise [192]. Pipelines [553]. Piranha [479]. Pivoting [158]. Places [481]. Planar [485, 313, 384, 59]. Plane [118, 113, 462, 548, 257, 596, 155, 11]. Planes [582]. Planetary [474, 479]. Planning [194]. Plans [245]. Plates [338]. Plick [143]. Point [211, 27, 525, 442]. Point-Free [211]. Points [544, 221, 13, 507]. Polar [341]. Polish [96]. Polygon [113, 114]. Polygons [349, 503]. Polyhedra [303]. Polymerase [362]. Polynomial [367, 61, 251, 594, 8, 160, 167, 161, 69, 597, 492, 456, 602, 203, 396, 206, 79, 508, 604, 68, 187, 138, 451, 589, 17, 137]. Polynomially [524]. Polynomials [409, 253, 428, 563, 19, 466, 483, 12, 18, 595, 523, 497, 183, 184, 120, 484, 565, 610, 16, 507, 437]. Polyominoes [554]. Polytope [422]. Polytopes [254, 494]. Pommaret [531]. Populations [65]. Possible [42]. Potential [383]. Potentials [615]. Power [195, 439]. Practical [103]. Practice [361]. Pre [218]. Pre- [218]. Precision [255, 397, 260]. Preemptive [79, 455]. Prefix [261]. Prescribed [606]. Preservation [165]. Primary [559]. Prime [489, 71, 559]. Principle [224]. Principled [169]. Prior [237]. Privacy [405]. Probabilities [26]. Problem [365, 496, 312, 377, 474, 178, 537, 384, 81, 587, 353, 159]. Problems [37, 272, 127, 333, 107, 457, 27, 603, 439, 263, 223, 317, 316, 319, 63, 336]. Procedure [204]. Process [372, 15, 444, 65, 263]. Processor [455]. Produce [487]. Producing [559]. Product [71, 227, 390, 151]. Products [342, 177, 227, 616]. Prof. [569]. Professor [403]. Profiles [362]. Program [495]. Programming [41]. Programs [495, 172]. Projection [401]. Projections [221]. Projective [179, 582]. Projects [94]. Proof [99, 50, 248]. Proofs [172, 224, 344, 276, 51]. Propagation [52]. Propelinear [417]. Proper [80]. Properties [142, 326, 539, 584, 111, 176, 391, 562, 497, 188, 503]. Property [165]. Protected [555]. Protein [362]. Protocols [187]. Prove [562]. Provers [319]. Provide [553]. Proving [315, 576, 432, 499]. Proximities [211, 391]. Proximity [215]. Pseudo [129]. Pseudo-Linear [129]. Pseudospectra [19]. Pseudozero [18]. Pseudozeros [17]. PTIME [176]. Publications [98]. Puiseux [519]. Pursuers [268]. Puzzles [557].

QE [302, 167]. QoS [207]. QoS-Aware [207]. Quadratic [431, 161, 396, 508]. Quadrilateral [577]. Quadrilaterals [589]. Qualification [270]. Qualitative [622]. Quality [468, 487]. Quantified [261]. Quantifier [261, 26]. Quantities [613]. Quantum [308, 280, 461]. Quasi [417, 211, 483, 437, 85]. Quasi-Cyclic [85]. Quasi-Hadamard [417]. Quasi-linear [483]. Quasi-Metrics [211]. Quasi-polynomials [437]. Quaternion [389]. Quaternionic [415]. Query [54]. Querying [47]. Questionnaire [350]. Queueing [444, 331]. Quiver [367]. Quotient [532].

Radical [21]. Radius [178]. Railway [245, 241, 452]. Ramsey [311]. Random [111, 25]. Randomized [106]. Range [74]. Rank [602, 22]. Rankin [379]. Ranking [55]. Rate [512]. Rates [377, 474]. Rational [258, 519, 113, 43, 201, 609, 617]. Rays [422]. Reachability [303, 266]. Reachable [267]. Reaction [61, 166, 228]. Ready [576]. Real [279, 118, 175, 466, 160, 18, 375, 280, 140, 273, 526, 197, 16, 550]. Real-Time [273, 197, 375]. Reasonable [176]. Reasoning [244, 46]. Recognition [97, 227]. Recognized [310]. Recognizing [520]. Recollections [569]. Reconsidered [616]. Reconstructing [228]. Rectilinear [224]. Recurrence [180]. Recurrences [577]. Recursive [300]. Recycling [248]. Reduced [68]. Reducibility [40]. Reduction [130, 61, 166, 373, 534]. Redundant [20, 86]. Refinement [125]. Reflections [449, 578, 585]. Reforms [514]. Regular [129, 123, 518, 429, 503, 328, 151]. Regularization [353]. Regulatory [165]. Related [147, 289]. Relations [43, 183, 356]. Relative [415, 373, 532]. Relator [348]. Relaxed [499, 203]. Relevance [55]. Reliability [406]. Remarks [244]. Repeats [103]. Representable [326].

Representation [323, 169, 46, 344]. Representations [360, 356, 472]. Represented [563]. Residues [441]. Resistance [433]. Resolvability [152]. Resonance [478, 594]. Resonances [612]. Resources [584]. Response [393]. Restrained [298]. Resultants [122, 14]. Results [27, 487, 83]. Retrieval [343, 214]. Revisited [180, 241]. Rewrite [101]. Rid [56]. Riemannian [568]. Riesz [353]. Right [382]. Ring [69, 68, 439, 108, 10]. Ringed [258]. Rings [367, 158, 532, 472, 68, 10]. Risk [605]. Ritt [346]. RMQ [107]. RMSQ [107]. RNA [31]. Robotics [194]. Robots [200]. Robust [422, 273, 274, 6]. Root [160, 67, 16]. Roots [181, 138]. Rota [131, 132]. Rotation [450]. Rows [557]. Rule [378]. Runge [445]. RWTH [570].

Safety [495, 176]. Sailboat [200]. SAT [110, 614, 106]. Satisfaction [272]. Satisfiability [304]. Satisfying [606]. Saving [14]. Scaffolding [136]. Scalability [104]. Scale [111]. Scales [534]. Scenarios [459]. Scheduling [455, 79]. Scheme [253, 245]. Schemes [613]. School [476]. Science [347, 276, 96]. Scientific [4, 174, 173, 98]. Scoring [35]. Scratch [401]. Sculptured [196]. Search [110, 590, 592, 77, 108, 55]. Searching [560]. Second [11, 507]. Second- [11]. Sections [382]. sectors [504]. Secure [498, 298]. Security [285, 404]. Segmentation [615]. Seidel [413]. Selected [586]. Self [146, 513, 542]. Self-organization [513]. Self-Paced [542]. Self-Stabilizing [146]. Semantic [229, 53, 345]. Semantics [191]. Semi [117, 444, 359, 32]. Semi-Algebraic [117]. Semi-external [359]. Semi-local [32]. Semicontinuity [566]. Semigroups [348]. Sensitivity [237]. Sensor [269]. Separable [617]. Separations [424]. Sequence [508, 362]. Sequences [355, 31, 421, 448]. Sequential [220]. Serbia [95]. Series [519, 307, 195, 439]. Serre [130]. Server [498]. Service [385, 277]. Set [54, 420, 18, 330, 266, 265]. Sets [285, 415, 323, 575, 332, 254, 214, 268, 117, 186, 215, 209, 210, 265, 267, 346, 218]. Shape [441]. Sheaf [218]. Shift [340]. Short [518]. Siamese [414]. Side [407]. Side-Channel [407]. Sign [8]. Signature [214, 469]. Signature-based [214]. Similar [355]. Similarities [548]. Simple [480, 4, 240]. Simplified [353]. Simulation [445, 516, 380]. Single [498, 281]. Singleton [59]. Singular [334, 340, 511, 133]. Singularities [526, 583]. Singularity [527]. SIRS [512]. Sixty [406]. Size [420, 105]. Skew [367, 482, 413]. Slabs [516]. Sliding [274]. Small [409, 89, 187]. Smooth [136, 602]. Smoothness [409]. SMS [570]. Snarks [295]. Snowflake [282]. Social [244]. Soft [233]. Software [494, 432, 476, 173]. Solid [273, 319, 587]. Solution [261, 245, 195, 607, 281]. Solutions [129, 519, 25, 606, 180, 201, 557]. Solve [63]. Solved [537]. Solvers [262]. Solving [614, 272, 89, 481, 365, 496, 603, 266, 399, 206, 439, 353, 159]. Some [244, 334, 483, 342, 449, 585, 13, 232, 405, 589, 461, 421, 475, 578]. Sorted [157]. Sorting [332]. SOS [167]. SOS-QE [167]. Source [446]. Space [33, 607, 14, 6, 305, 353]. Spaced [358]. Spaces [345, 189, 252, 215, 422, 250, 309, 216]. Spain [94]. Spanning [73]. Special [92, 235, 500, 511, 301, 160, 193, 335, 392, 506, 325, 276]. Specialization [495]. Specific [245, 553]. Specification [239]. Specifying [101]. Spectra [334]. Spectral [229]. Spiders [310]. Spiric [600]. Spline [252, 137, 222]. Splines [5, 487]. Spread [150, 543]. Spring [478]. Spurious [56]. Squares [485]. Stability [512, 612, 64, 164, 322, 28]. Stabilizing [146]. Standard [566]. Star [227, 588, 311].

State [144, 385, 520, 200, 338, 274].

Stationary [560]. Statistical [404]. Steady [520]. STEM [543]. Step [161, 281]. Stochastic [62, 171, 26]. Stokes [515]. Strategies [101, 262]. Strength [242, 483, 289, 288, 297]. Stress [338]. String [355, 363, 365, 103, 32]. Strings [356]. Strong [71, 294, 293]. Structural [326, 553]. Structure [618, 374, 568, 362]. Structured [22]. Structures [24, 598, 541, 556, 352, 551]. Students [430, 376, 475]. Studies [248, 108]. Study [340, 511, 369, 429, 379, 389, 608]. Sub [256]. Sub-Decompositions [256]. Subclass [451]. Subdivision [262]. Subdivisions [599]. Submanifolds [136]. Subnetwork [60]. Subpavings [267]. Subscripts [49]. Subsequence [364]. Subset [349]. Substitution [220]. Substrings [30]. Subway [471]. Success [449, 585]. Successive [190]. Succinct [360]. Suffix [358, 357]. Suitable [599]. Suite [174]. Sum [349]. Summands [240]. Summation [240]. Summations [448]. Sums [533, 150]. Super [226, 149]. Superelliptic [91]. Superposition [204, 205]. Superquadratic [161]. Superscripts [49]. Superspecial [590, 592]. Supplemented [432]. Support [411, 378]. Supra [142]. Supra-Hereditary [142]. Surface [88, 195, 196, 197]. Surfaces [258, 115, 116, 389]. Survey [343, 177, 597, 404]. SVD [21]. Swarm [313]. Sweep [281]. Swinging [380]. Swiss [159]. swMATH [385]. Symbol [601, 11]. Symbolic [529, 124, 581, 334, 340, 511, 379, 530, 168, 335, 392, 506, 396, 65, 604, 14, 502, 608]. Symbolic-Numeric [529, 168]. Symbolically [606]. Symmetric [413]. Symmetries [124, 190]. Symmetry [457, 524, 324]. Symmetry-Based [524]. Synchronization [78, 513]. Syntactic [598, 541, 556, 352, 551]. Syntax [387]. Synthesizing [196]. Synthetic [536].

System [495, 530, 560, 432, 486, 447, 479, 196, 449, 585, 383, 622, 505, 502, 26]. Systems [202, 129, 612, 130, 61, 166, 89, 519, 483, 535, 302, 25, 303, 351, 201, 368, 167, 523, 161, 246, 27, 492, 322, 203, 206, 247, 439, 28, 273, 578, 583, 51]. Syzygies [456].

Tactics [99]. Tale [600]. Tamari [574]. Task [455]. TASS [173]. Tau [333, 603, 399, 398]. Teaching [371, 542]. Techniques [460, 386, 405, 274, 32]. Technologies [580]. Telescoping [128, 533]. Temporal [46]. Ternary [356]. Test [409, 186]. Testing [466, 8, 546, 545, 525]. Tests [4]. Tetrahedral [294]. Text [219, 363]. th [181]. Their [310, 548, 535, 551, 405]. Them [56]. Theorem [491, 499, 319]. Theoretic [218]. Theories [205]. Theory [202, 133, 518, 244, 581, 430, 212, 200, 308, 479, 325, 383]. Thermal [516]. Thinning [20]. Third [604, 11]. Third- [11]. Three [348, 338, 377, 474, 297]. Three-Body [377, 474]. Three-Dimensional [338]. Through-Cracked [338]. TiGL [446]. Tikhonov [353]. Tilted [589]. Time [515, 272, 33, 535, 534, 224, 330, 79, 47, 108, 273, 449, 197, 375]. Time-Dependent [272]. Time-Fractional [535]. Time-Lines [47]. Time-Optimal [224]. Timed [204]. Tolerant [288]. Tool [432, 195]. Toolbox [603]. Toolkit [336, 173]. Tools [442]. Topological [425, 332, 556]. Topology [552, 118, 117, 387]. Tori [600, 291]. Toric [593]. Toricity [520]. Toroidal [242]. Torsion [527]. Torus [290]. Total [242, 289, 288, 297]. Traceability [145]. Traces [88]. Tracking [274]. Tractable [38]. Trajectory [194, 274]. Transcendental [181]. Transform [324, 359]. Transformation [52]. Transformations [363]. Transforming [98]. Transforms [536]. Transition [338]. Transitions [156]. Transitive [579].

Translational [389]. TransLucid [41]. Transmission [584, 337]. Transmit [584]. Transparent [378]. Tree [146, 97]. Trees [73, 111, 360, 148, 155]. Triangle [318, 316, 504]. Triangular [523]. Triangulating [136]. Triangulation [305]. Triangulations [468]. Trifolium [429]. Trivariate [201, 252]. Tropical [494, 522]. Tropicalization [250]. Truncation [184]. Truth [315]. Tuberculosis [513]. Turtle [543]. Tutorial [36]. Twin [414]. Two [548, 600, 509, 530, 596, 606, 368, 188, 550, 268, 224, 602, 312, 474, 181, 85, 299, 125]. Two-Connected [299]. Two-Dimensional [602]. Two-Edge [299]. Two-Factor [125]. Two-Parameter [509]. Two-phase [606]. Two-Planetary [474]. Type [230, 565, 140, 611, 594, 120]. Types [232].

Ugly [93]. Ultrafilter [216]. Unambiguous [34]. Unavoidable [315]. Uncertainty [315]. Underwater [271]. Uniform [292]. Uniformities [212]. Union [226]. Unistochastic [422]. Unitals [582]. univalent [451]. Univariate [16]. Universal [321]. Unnecessary [559]. Unusual [35]. Unweighted [178]. Upper [521]. Usage [386]. Use [430, 442]. Used [584]. User [279]. User-Friendly [279]. Using [614, 401, 255, 33, 269, 462, 485, 566, 15, 596, 486, 454, 21, 434, 53, 375, 164, 190, 46, 241, 543, 271, 622, 487, 502, 389, 376, 475, 166, 213]. Utility [57].

Validated [175]. Validation [529]. Value [268, 605, 383]. Value-at-Risk [605]. Variable [377, 304, 281]. Variables [89, 181]. Variants [37, 89]. Variational [222]. Varieties [520, 525]. Varying [35, 474]. Vehicles [313, 271]. Verhulst [25]. Verification [172, 176, 97, 262, 247, 452, 260, 174, 51]. Verification-Driven [247]. Verifying [245]. Version [77, 318, 344]. Versioning [42].

Versus [266]. Vertex [288, 154, 459]. Vertices [305]. Via [439, 324, 495, 593, 220, 12, 419, 492, 508, 604, 551, 583]. VIBes [275]. Viral [362]. Virtual [515, 96]. Virus [543]. Viscous [281]. Visibility [155]. Visualisation [477]. Visualizer [275]. Visually [543]. Vladimir [569, 572, 528, 571, 570, 567]. Volterra [368]. Volumes [109]. Voronoï [552].

Wadge [43, 39, 40]. way [400]. Weak [410]. Web [289]. Weber [517]. Weighted [37, 499, 437]. Weil [558]. well [274]. Weyl [367]. Wheeler [359]. Wheels [230]. Whose [305]. Wildcards [54]. Wirelength [153]. Wireless [586]. Within [398, 245]. Without [370, 559]. Witness [254]. Word [326, 105]. Words [35, 601, 131, 132, 544]. Worlds [42].

XaoS [549]. XML [98]. XOR [614].

Year [449]. Years [347, 406, 571]. Young [372].

Zero [69, 179, 525, 4]. Zero-Dimensionality [525]. Zeros [523, 6]. Zeta [91]. Ziv [33]. Zonal [497]. Zonotope [266]. Zonotope-Based [266].

References

Wang:2007:E

[1] Dongming Wang and Zhiming Zheng. Editorial. Mathematics in Computer Science, 1(1):1–2, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=1&spage=1.

Yap:2007:F

[2] Chee K. Yap and Hoon Hong. Foreword. Mathematics in Computer Science, 1(1):3–7, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=1&spage=3.

Trefethen:2007:CNF

[3] Lloyd N. Trefethen. Computing numerically with functions instead of numbers. Mathematics in Computer Science, 1(1):9–19, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=1&spage=9.

Richardson:2007:ZTC

[4] Daniel Richardson. Zero tests for constants in simple scientific computation. Mathematics in Computer Science, 1(1):21–37, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=1&spage=21.

Ghosh:2007:ACS

[5] Sunayana Ghosh, Sylvain Petitjean, and Gert Vegter. Approximation by conic splines. Mathematics in Computer Science, 1(1):39–69, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=1&spage=39.

Sharma:2007:RAZ

[6] Vikram Sharma. Robust approximate zeros in Banach space. Mathematics in Computer Science, 1(1):71–109, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=1&spage=
71.

vanderHoeven:2007:EAC

[7] Joris van der Hoeven. On effective analytic continuation. Mathematics in Computer Science, 1(1):111–175, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=1&spage=
111.

ElDin:2007:TSC

[8] Mohab Safey El Din. Testing sign conditions on a multivariate polynomial and applications. Mathematics in Computer Science, 1(1):177–207, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=1&spage=177.

Wang:2007:F

[9] Dongming Wang and Lihong Zhi. Foreword. Mathematics in Computer Science, 1(2):209–210, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=2&spage=209.

Zhou:2007:CGB

[10] Meng Zhou and Franz Winkler. On computing Gröbner bases in rings of differential operators with coefficients in a ring. Mathematics in Computer Science, 1(2):211–223, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=2&spage=211.

Shemyakova:2007:PFS

[11] Ekaterina Shemyakova. Parametric factorizations of second-, third- and Fourth–Order linear partial differential operators with a completely factorable symbol on the plane. Mathematics in Computer Science, 1(2):225–237, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=2&spage=
225.

Feinsilver:2007:IAF

[12] Philip Feinsilver and René Schott. Inversion of analytic functions via canonical polynomials: a matrix approach. Mathematics in Computer Science, 1 (2):239–251, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=2&spage=239.

Romanovski:2007:BPP

[13] Valery G. Romanovski. Bifurcations of periodic points of some algebraic maps. Mathematics in Computer Science, 1(2):253–265, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=2&spage=253.

Pearson:2007:SSC

[14] Jane M. Pearson and Noel G. Lloyd. Space saving calculation of symbolic resultants. Mathematics in Computer Science, 1(2):267–290, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=2&spage=267.

Gu:2007:UCA

[15] Nong Gu, Daniel Lazard, Fabrice Rouillier, and Yong Xiang. Using computer algebra to certify the global convergence of a numerical optimization process. Mathematics in Computer Science, 1(2):291–304, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=2&spage=291.

Zhang:2007:NMR

[16] Ting Zhang and Bican Xia. A new method for real root isolation of univariate polynomials. Mathematics in Computer Science, 1(2):305–320, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=2&spage=
305.

Sekigawa:2007:LPC

[17] Hiroshi Sekigawa and Kiyoshi Shirayanagi. On the location of pseudozeros of a complex interval polynomial. Mathematics in Computer Science, 1(2):321–335, December 2007. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=1&issue=2&spage=321.

Graillat:2007:PSR

[18] Stef Graillat. Pseudozero set of real multivariate polynomials. Mathematics in Computer Science, 1(2):337–352, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=1&issue=2&spage=
                                        337.

Corless:2007:PMP

[19] Robert M. Corless and Nargol Rezvani. Pseudospectra of matrix polynomials that are expressed in alternative bases. Mathematics in Computer Science, 1(2):353–374, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-007-0010-x.

Abbott:2007:TRE

[20] John Abbott, Claudia Fassino, and Maria-Laura Torrente. Thinning out redundant empirical data. Mathematics in Computer Science, 1(2):375– 392, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s11786-007-
0020-8.

Janovitz-Freireich:2007:ARC

[21] Itnuit Janovitz-Freireich, Lajos Rónyai, and Ágnes Szántó. Approximate radical for clusters: A global approach using Gaussian elimination or SVD. Mathematics in Computer Science, 1(2):393–425, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-007-0013-7.

Sun:2007:SLR

[22] Dongxia Sun and Lihong Zhi. Structured low rank approximation of a Bezout matrix. Mathematics in Computer Science, 1(2):427–437, December 2007. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-007-0014-6.

Kotsireas:2008:F

[23] Ilias S. Kotsireas. Foreword. Mathematics in Computer Science, 1(3):439, March 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=3&spage=
439.

Chen:2008:CSN

[24] William Y. C. Chen, Andreas W. M. Dress, and Winking Q. Yu. Community structures of networks. Mathematics in Computer Science, 1(3):441–457, March 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1661-8270&volume=1&issue=3&spage=
                                        441.

Ganzha:2008:ESH

[25] Elena I. Ganzha, Valery M. Loginov, and Sergey P. Tsarev. Exact solutions of hyperbolic systems of kinetic equations. application to Verhulst model with random perturbation. Mathematics in Computer Science, 1(3):459–472, March 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=1&issue=3&spage=
                                        459.

Yoshida:2008:IPS

[26] Hiroshi Yoshida, Katsuhisa Horimoto, and Hirokazu Anai. Inference of probabilities over a stochastic IL-system by quantifier elimination. Mathematics in Computer Science, 1(3):473–485, March 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=1&issue=3&spage=
                                        473.

Kosub:2008:DRF

[27] Sven Kosub. Dichotomy results for fixed-point existence problems for Boolean dynamical systems. Mathematics in Computer Science, 1(3):487–505, March 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=1&issue=3&spage=
                                        487.

Niu:2008:AAS

[28] Wei Niu and Dongming Wang. Algebraic approaches to stability analysis of biological systems. Mathematics in Computer Science, 1(3):507–539, March 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=3&spage=
507.

Chan:2008:F

[29] Joseph Wun-Tat Chan and Maxime Crochemore. Foreword. Mathematics in Computer Science, 1(4):541–542, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=4&spage=
541.

Chan:2008:NOC

[30] H. L. Chan, T. W. Lam, W. K. Sung, P. W. H. Wong, and S. M. Yiu. Nonoverlapping common substrings allowing mutations. Mathematics in Computer Science, 1(4):543–555, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=4&spage=543.

Iliopoulos:2008:NAP

[31] Costas S. Iliopoulos, Laurent Mouchard, and M. Sohel Rahman. A new approach to pattern matching in degenerate DNA/RNA sequences and distributed pattern matching. Mathematics in Computer Science, 1(4):557–569, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=4&spage=
557.

Tiskin:2008:SLS

[32] Alexander Tiskin. Semi-local string comparison: Algorithmic techniques and applications. Mathematics in Computer Science, 1(4):571–603, June 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=4&spage=571.

Chen:2008:LZF

[33] Gang Chen, Simon J. Puglisi, and W. F. Smyth. Lempel–Ziv factorization using less time & space. Mathematics in Computer Science, 1(4):605–623, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=4&spage=605.

Beal:2008:UA

[34] Marie-Pierre Béal, Eugen Czeizler, Jarkko Kari, and Dominique Perrin. Unambiguous automata. Mathematics in Computer Science, 1(4):625–638, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=1&issue=4&spage=
625.

Apostolico:2008:SUW

[35] Alberto Apostolico and Cinzia Pizzi. Scoring unusual words with varying mismatch errors. Mathematics in Computer Science, 1(4):639–653, June 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=4&spage=639.

Giancarlo:2008:TCC

[36] Raffaele Giancarlo, Davide Scaturro, and Filippo Utro. A tutorial on computational cluster analysis with applications to pattern discovery in microarray data. Mathematics in Computer Science, 1(4):655–672, June 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=4&spage=655.

Banerjee:2008:EAV

[37] Satyajit Banerjee, Atish Datta Chowdhury, and Subhas Kumar Ghosh. Efficient algorithms for variants of weighted matching and assignment problems. Mathematics in Computer Science, 1 (4):673–688, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=1&issue=4&spage=673.

Farago:2008:GTD

[38] András Faragó. A general tractable density concept for graphs. Mathematics in Computer Science, 1(4):689–699, June 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1661-8270&volume=1&issue=4&spage=
                                        689.

Plaice:2008:HBW

[39] John Plaice and Mehmet A. Orgun. In honour of Bill Wadge. Mathematics in Computer Science, 2(1):1–3, November 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=2&issue=1&spage=
                                        1.

Selivanov:2008:WRI

[40] Victor Selivanov. Wadge reducibility and infinite computations. Mathematics in Computer Science, 2(1):5–36, November 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=2&issue=1&spage=
                                        5.

Plaice:2008:LTI

[41] John Plaice, Blanca Mancilla, and Gabriel Ditu. From Lucid to TransLucid: Iteration, dataflow, intensional and Cartesian programming. Mathematics in Computer Science, 2(1): 37–61, November 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=2&issue=1&spage=37.

Mancilla:2008:PWV

[42] Blanca Mancilla and John Plaice. Possible worlds versioning. Mathematics in Computer Science, 2(1):63–83, November 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=1&spage=
63.

Finkel:2008:WDI

[43] Olivier Finkel. Wadge degrees of infinitary rational relations. Mathematics in Computer Science, 2(1):85–102, November 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=1&spage=
85.

Orchard:2008:ILD

[44] Dominic A. Orchard and Steve Matthews. Integrating Lucid’s declarative dataflow paradigm into object-orientation. Mathematics in Computer Science, 2(1): 103–122, November 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=1&spage=103.

Charalambidis:2008:EII

[45] Angelos Charalambidis, Athanasios Grivas, Nikolaos S. Papaspyrou, and Panos Rondogiannis. Efficient intensional implementation for lazy functional languages. Mathematics in Computer Science, 2(1):123–141, November 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=1&spage=123.

Orgun:2008:KRR

[46] Mehmet A. Orgun, Chuchang Liu, and Abhaya C. Nayak. Knowledge representation, reasoning and integration using temporal logic with clocks. Mathematics in Computer Science, 2 (1):143–163, November 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=1&spage=143.

Orgun:2008:QHD

[47] Mehmet A. Orgun. Querying historical data over multiple time-lines. Mathematics in Computer Science, 2 (1):165–191, November 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=1&spage=165.

Kerber:2008:F

[48] Manfred Kerber. Foreword. Mathematics in Computer Science, 2(2): 193–194, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=2&spage=193.

Aly:2008:ISS

[49] Walaa Aly, Seiichi Uchida, and Masakazu Suzuki. Identifying subscripts and superscripts in mathematical documents. Mathematics in Computer Science, 2(2):195–209, December 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=2&spage=195.

Cohen:2008:APG

[50] Arjeh M. Cohen, Jan Willem Knopper, and Scott H. Murray. Automatic proof of graph nonisomorphism. Mathematics in Computer Science, 2(2):211–229, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=2&spage=
211.

Urban:2008:ABC

[51] Josef Urban and Geoff Sutcliffe. ATPbased Cross–Verification of Mizar proofs: Method, systems, and first experiments. Mathematics in Computer Science, 2(2):231–251, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=2&spage=231.

Autexier:2008:OTP

[52] Serge Autexier, Christoph Benzmüller, Dominik Dietrich, and Marc Wagner. Organization, transformation, and propagation of mathematical knowledge in Ω mega. Mathematics in Computer Science, 2(2):253–277, December 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=2&spage=253.

Kohlhase:2008:ULS

[53] Michael Kohlhase. Using LaTEX as a semantic markup format. Mathematics in Computer Science, 2(2):279–304, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=2&issue=2&spage=
                                        279.

Altamimi:2008:MQL

[54] Moody Ebrahem Altamimi and Abdou Youssef. A math query language with an expanded set of wildcards. Mathematics in Computer Science, 2(2):305–331, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=2&issue=2&spage=
                                        305.

Youssef:2008:RRH

[55] Abdou S. Youssef. Relevance ranking and hit description in math search. Mathematics in Computer Science, 2 (2):333–353, December 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=2&issue=2&spage=333.

Coen:2008:SDE

[56] Claudio Sacerdoti Coen and Stefano Zacchiroli. Spurious disambiguation errors and how to get rid of them. Mathematics in Computer Science, 2(2):355–378, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=2&issue=2&spage=
                                        355.

Davenport:2008:FEO

[57] James H. Davenport and Paul Libbrecht. The freedom to extend Open-Math and its utility. Mathematics in Computer Science, 2(2):379–398, December 2008. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=2&spage=
379.

Anai:2009:F

[58] Hirokazu Anai and Katsuhisa Horimoto. Foreword. Mathematics in Computer Science, 2(3):399–400, March 2009. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=3&spage=399.

Tamura:2009:ASA

[59] Takeyuki Tamura and Tatsuya Akutsu. Algorithms for singleton attractor detection in planar and nonplanar AND/ OR Boolean networks. Mathematics in Computer Science, 2(3):401–420, March 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=3&spage=
401.

Siebert:2009:DBB

[60] Heike Siebert. Deriving behavior of Boolean bioregulatory networks from subnetwork dynamics. Mathematics in Computer Science, 2(3):421–442, March 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=3&spage=

Boulier:2009:TAR

[61] François Boulier, François Lemaire, Alexandre Sedoglavic, and Asli Ürgüplü. Towards an automated reduction method for polynomial ODE models of biochemical reaction systems. Mathematics in Computer Science, 2 (3):443–464, March 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=2&issue=3&spage=443.

Bortolussi:2009:HDS

[62] Luca Bortolussi and Alberto Policriti. Hybrid dynamics of stochastic πcalculus. Mathematics in Computer Science, 2(3):465–491, March 2009. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=2&issue=3&spage=465.

Sturm:2009:IAL

[63] Thomas Sturm, Andreas Weber, Essam O. Abdel-Rahman, and M’hammed El Kahoui. Investigating algebraic and logical algorithms to solve Hopf bifurcation problems in algebraic biology. Mathematics in Computer Science, 2(3):493–515, March 2009. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=2&issue=3&spage=493.

ElKahoui:2009:SDF

[64] Mh́ammed El Kahoui and Adamou Otto. Stability of disease free equilibria in epidemiological models. Mathematics in Computer Science, 2(3):517–533, March 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=3&spage=
517.

McCaig:2009:IPS

[65] Chris McCaig, Rachel Norman, and Carron Shankland. From individuals to populations: a symbolic process algebra approach to epidemiology. Mathematics in Computer Science, 2(3):535–556, March 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=3&spage=
535.

Ilie:2009:CCE

[66] Silvana Ilie, Robert M. Corless, and Chris Essex. The computational complexity of extrapolation methods. Mathematics in Computer Science, 2 (4):557–566, December 2009. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=4&spage=557.

Pikkarainen:2009:BMR

[67] Hanna K. Pikkarainen and Josef Schicho. A Bayesian model for root computation. Mathematics in Computer Science, 2(4):567–586, December 2009. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=2&issue=4&spage=567.

Nabeshima:2009:RGB

[68] Katsusuke Nabeshima. Reduced Gröbner bases in polynomial rings over a polynomial ring. Mathematics in Computer Science, 2(4):587–599, December 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=2&issue=4&spage=
                                        587.

Kapur:2009:ACG

[69] Deepak Kapur and Yongyang Cai. An algorithm for computing a Gröbner basis of a polynomial ideal over a ring with zero divisors. Mathematics in Computer Science, 2(4):601–634, December 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=2&issue=4&spage=601.

Alcazar:2009:GLB

[70] Juan Gerardo Alcázar. Good local behavior of offsets to implicit algebraic curves. Mathematics in Computer Science, 2(4):635–652, December 2009. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=2&issue=4&spage=635.

Hellmuth:2009:LAP

[71] Marc Hellmuth, Wilfried Imrich, Werner Klöckl, and Peter F. Stadler. Local algorithms for the prime factorization of strong product graphs. Mathematics in Computer Science, 2(4):653–682, December 2009. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=2&issue=4&spage=
653.

Miller:2010:F

[72] Mirka Miller and Koichi Wada. Foreword. Mathematics in Computer Science, 3(1):1–2, March 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=1.

Baumgart:2010:PBG

[73] Matthias Baumgart. Partitioning bispanning graphs into spanning trees. Mathematics in Computer Science, 3(1):3–15, March 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=3.

Fischer:2010:FRM

[74] Johannes Fischer and Volker Heun. Finding range minima in the middle: Approximations and applications. Mathematics in Computer Science, 3(1):17–30, March 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=17.

Fujita:2010:LCG

[75] Satoshi Fujita. Loose cover of graphs. Mathematics in Computer Science, 3 (1):31–38, March 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=31.

Ichishima:2010:PLG

[76] Rikio Ichishima and Akito Oshima. On partitional labelings of graphs. Mathematics in Computer Science, 3(1):39–45, March 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=3&issue=1&spage=
39.

Kutzelnigg:2010:IVC

[77] Reinhard Kutzelnigg. An improved version of cuckoo hashing: Average case analysis of construction cost and search operations. Mathematics in Computer Science, 3(1):47–60, March 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=47.

Lee:2010:POS

[78] Sof Anthony Lee. k-phase oscillator synchronization for graph coloring. Mathematics in Computer Science, 3(1):61–72, March 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=3&issue=1&spage=
61.

Mertzios:2010:PSE

[79] George B. Mertzios and Walter Unger. Preemptive scheduling of equal-length jobs in polynomial time. Mathematics in Computer Science, 3(1):73–84, March 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=3&issue=1&spage=
73.

Mertzios:2010:OAF

[80] George B. Mertzios and Walter Unger. An optimal algorithm for the k-fixed-endpoint path cover on proper interval graphs. Mathematics in Computer Science, 3(1):85–96, March 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=85.

Sivan:2010:CCP

[81] Balasubramanian Sivan, S. Harini, and C. Pandu Rangan. On conditional covering problem. Mathematics in Computer Science, 3(1):97–107, March 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=97.

Tanaka:2010:PCC

[82] Yuuki Tanaka and Yukio Shibata. On the pagenumber of the cube-connected cycles. Mathematics in Computer Science, 3(1):109–117, March 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=1&spage=109.

Tang:2010:NRE

[83] Jianmin Tang, Yuqing Lin, and Mirka Miller. New results on EX graphs.

Mathematics in Computer Science, 3 (1):119–126, March 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=1&spage=119.

Faugere:2010:F

[84] Jean-Charles Faugère and Ludovic Perret. Foreword. Mathematics in Computer Science, 3(2):127–128, April 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=2&spage=127.

Otmani:2010:CTM

[85] Ayoub Otmani, Jean-Pierre Tillich, and Léonard Dallot. Cryptanalysis of two McEliece cryptosystems based on quasicyclic codes. Mathematics in Computer Science, 3(2):129–140, April 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=2&spage=129.

Heuberger:2010:RAE

[86] Clemens Heuberger. Redundant τadic expansions II: Non-optimality and chaotic behaviour. Mathematics in Computer Science, 3(2):141–157, April 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=3&issue=2&spage=
                                        141.

Eibach:2010:OGB

[87] Tobias Eibach, Gunnar Völkel, and Enrico Pilz. Optimising Gröbner bases on Bivium. Mathematics in Computer Science, 3(2):159–172, April 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=2&spage=159.

Duquesne:2010:TGL

[88] Sylvain Duquesne. Traces of the group law on the Kummer surface of a curve of genus 2 in characteristic 2. Mathematics in Computer Science, 3(2):173–183, April 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=3&issue=2&spage=
173.

Bulygin:2010:OSS

[89] Stanislav Bulygin and Michael Brickenstein. Obtaining and solving systems of equations in key variables only for the small variants of AES. Mathematics in Computer Science, 3(2):185–200, April 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=3&issue=2&spage=
185.

Lee:2010:CGC

[90] Moon Sung Lee and Sang Geun Hahn. Cryptanalysis of the GGH cryptosystem. Mathematics in Computer Science, 3(2):201–208, April 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=2&spage=201.

Minzlaff:2010:CZF

[91] Moritz Minzlaff. Computing zeta functions of superelliptic curves in larger characteristic. Mathematics in Computer Science, 3(2):209–224, April 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=2&spage=209.

Autexier:2010:FSI

[92] Serge Autexier, Petr Sojka, and Masakazu Suzuki. Foreword to the special issue on authoring, digitalization and management of mathematical knowledge. Mathematics in Computer Science, 3(3):225–226, May 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=3&spage=225.

Bouche:2010:DML

[93] Thierry Bouche. Digital mathematics libraries: The good, the bad, the ugly. Mathematics in Computer Science, 3(3):227–241, May 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=3&spage=227.

Macias-Virgos:2010:DPS

[94] E. Macías-Virgós and R. de la Viesca. Digitization projects in Spain.

Mathematics in Computer Science, 3 (3):243–250, May 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=3&spage=243.

Mijajlovic:2010:DME

[95] Zarko Mijajlovič, Zoran Ognjanovič, and Aleksandar Pejovič. Digitization of mathematical editions in Serbia. Mathematics in Computer Science, 3(3):251–263, May 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=3&spage=251.

Zamlynska:2010:EMC

[96] Katarzyna Zamlyńska, Alek Tarkowski, and Tomasz Rosiek. Evolution of the mathematical collection of the Polish Virtual Library of Science. Mathematics in Computer Science, 3(3):265–278, May 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=3&issue=3&spage=
265.

Fujiyoshi:2010:GVM

[97] Akio Fujiyoshi, Masakazu Suzuki, and Seiichi Uchida. Grammatical verification for mathematical formula recognition based on context-free tree grammar. Mathematics in Computer Science, 3(3):279–298, May 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=3&spage=279.

Stamerjohanns:2010:TLC

[98] Heinrich Stamerjohanns, Michael Kohlhase, Deyan Ginev, Catalin David, and Bruce Miller. Transforming large collections of scientific publications to XML. Mathematics in Computer Science, 3(3):299–307, May 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=3&spage=299.

Aspinall:2010:THP

[99] David Aspinall, Ewen Denney, and Christoph Lüth. Tactics for hierarchical proof. Mathematics in Computer Science, 3(3):309–330, May 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=3&spage=309.

Cohen:2010:MCI

[100] A. M. Cohen, H. Cuypers, and R. Verrijzer. Mathematical context in interactive documents. Mathematics in Computer Science, 3(3):331–347, May 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=3&issue=3&spage=331.

Heeren:2010:SRS

[101] Bastiaan Heeren, Johan Jeuring, and Alex Gerdes. Specifying rewrite strategies for interactive exercises. Mathematics in Computer Science, 3(3):349–370, May 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=3&issue=3&spage=
349.

Christodoulakis:2010:F

[102] Manolis Christodoulakis and Costas S.

Iliopoulos. Foreword. Mathematics in Computer Science, 3(4):371, June 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=371.

Puglisi:2010:FPA

[103] Simon J. Puglisi, W. F. Smyth, and Munina Yusufu. Fast, practical algorithms for computing all the repeats in a string. Mathematics in Computer Science, 3(4):373–389, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=373.

Cinque:2010:SCP

[104] Luigi Cinque, Sergio De Agostino, and Luca Lombardi. Scalability and communication in parallel low-complexity lossless compression. Mathematics in Computer Science, 3(4):391–406, June 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=391.

Kulekci:2010:BNB

[105] M. Oguzhan Külekci. BLIM: a new bit-parallel pattern matching algorithm overcoming computer word size limitation. Mathematics in Computer Science, 3(4):407–420, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=407.

Ghosh:2010:RAS

[106] Subhas Kumar Ghosh and Janardan Misra. A randomized algorithm for 3-SAT. Mathematics in Computer Science, 3(4):421–431, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=421.

Hasan:2010:COA

[107] Masud Hasan, Tanaeem M. Moosa, and M. Sohel Rahman. Cache oblivious algorithms for the RMQ and the RMSQ problems. Mathematics in Computer Science, 3(4):433–442, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=433.

Pu:2010:ASE

[108] Ida M. Pu and Yuji Shen. Analytical studies of energy-time efficiency of blocking expanding ring search. Mathematics in Computer Science, 3(4):443–456, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=443.

Dimitrova:2010:EVC

[109] Elena S. Dimitrova. Estimating the volumes of the cones in a Gröbner fan. Mathematics in Computer Science, 3(4):457–463, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=457.

Albrecht:2010:ALS

[110] A. A. Albrecht, P. C. R. Lane, and K. Steinhöfel. Analysis of local search landscapes for k-SAT instances. Mathematics in Computer Science, 3(4):465–488, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=465.

Cooper:2010:SFP

[111] Colin Cooper and Ryuhei Uehara.

Scale free properties of random ktrees. Mathematics in Computer Science, 3(4):489–496, June 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=3&issue=4&spage=489.

Gonzalez-Vega:2010:F

[112] Laureano Gonzalez-Vega and Sylvain Lazard. Foreword. Mathematics in Computer Science, 4(1):1–2, November 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1661-8270&volume=4&issue=1&spage=
                                        1.

DAndrea:2010:NPR

[113] Carlos D’Andrea and Martín Sombra. The Newton polygon of a rational plane curve. Mathematics in Computer Science, 4(1):3–24, November 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=1&spage=3.

Emiris:2010:CNP

[114] Ioannis Z. Emiris, Christos Konaxis, and Leonidas Palios. Computing the Newton polygon of the implicit equation. Mathematics in Computer Science, 4(1):25–44, November 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=1&spage=25.

Berberich:2010:APSa

[115] Eric Berberich, Efi Fogel, Dan Halperin, Kurt Mehlhorn, and Ron Wein. Arrangements on parametric surfaces I: General framework and infrastructure. Mathematics in Computer Science, 4(1):45–66, November 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=1&spage=45.

Berberich:2010:APSb

[116] Eric Berberich, Efi Fogel, Dan Halperin, Michael Kerber, and Ophir Setter. Arrangements on parametric surfaces II: Concretizations and applications. Mathematics in Computer Science, 4 (1):67–91, November 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=1&spage=67.

Lazard:2010:CTS

[117] Daniel Lazard. CAD and topology of semi-algebraic sets. Mathematics in Computer Science, 4(1):93–112, November 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=4&issue=1&spage=
93.

Cheng:2010:TRA

[118] Jinsan Cheng, Sylvain Lazard, Luis Peñaranda, Marc Pouget, Fabrice Rouillier, et al. On the topology of real algebraic plane curves. Mathematics in Computer Science, 4(1): 113–137, November 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=1&spage=113.

Regensburger:2010:F

[119] Georg Regensburger, Markus Rosenkranz, and William Y. Sit. Foreword. Mathematics in Computer Science, 4(2– 3):139–141, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=2&spage=139.

Levin:2010:DPI

[120] Alexander Levin. Dimension polynomials of intermediate fields and krulltype dimension of finitely generated differential field extensions. Mathematics in Computer Science, 4(2– 3):143–150, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=2&spage=143.

Morikawa:2010:DBE

[121] Shuji Morikawa, Katsunori Saito, Taihei Takeuchi, and Hiroshi Umemura. Discrete Burgers’ equation, binomial coefficients and mandala. Mathematics in Computer Science, 4(2– 3):151–167, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=2&spage=151.

Kasman:2010:FRP

[122] Alex Kasman and Emma Previato. Factorization and resultants of partial differential operators. Mathematics in Computer Science, 4(2– 3):169–184, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=2&spage=169.

Boulier:2010:NFA

[123] François Boulier and François Lemaire. A normal form algorithm for regular differential chains. Mathematics in Computer Science, 4(2–3):185–201, September 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=4&issue=2&spage=
185.

Cheviakov:2010:SCL

[124] Alexei F. Cheviakov. Symbolic computation of local symmetries of nonlinear and linear partial and ordinary differential equations. Mathematics in Computer Science, 4(2–3):203–222, September 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=4&issue=2&spage=
203.

Shemyakova:2010:RTF

[125] Ekaterina Shemyakova. Refinement of two-factor factorizations of a linear partial differential operator of arbitrary order and dimension. Mathematics in Computer Science, 4(2– 3):223–230, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=2&spage=223.

Plesken:2010:LDE

[126] Wilhelm Plesken and Daniel Robertz.

Linear differential elimination for analytic functions. Mathematics in Computer Science, 4(2–3):231–242, September 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1661-8270&volume=4&issue=2&spage=
                                        231.

Dimovski:2010:OCA

[127] Ivan Dimovski and Margarita Spiridonova. Operational calculus approach to nonlocal Cauchy problems. Mathematics in Computer Science, 4(2– 3):243–258, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=2&spage=243.

Koutschan:2010:FAC

[128] Christoph Koutschan. A fast approach to creative telescoping. Mathematics in Computer Science, 4(2– 3):259–266, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=2&spage=259.

Barkatou:2010:MMM

[129] Moulay A. Barkatou, Gary Broughton, and Eckhard Pflügel. A monomial-by-monomial method for computing regular solutions of systems of pseudolinear equations. Mathematics in Computer Science, 4(2–3):267–288, September 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=4&issue=2&spage=
                                        267.

Boudellioua:2010:SRL

[130] M. S. Boudellioua and A. Quadrat. Serre’s reduction of linear functional systems. Mathematics in Computer Science, 4(2–3):289–312, September 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=2&spage=289.

Guo:2010:EGFa

[131] Li Guo and William Y. Sit. Enumeration and generating functions of Rota–Baxter words. Mathematics in Computer Science, 4(2–3):313–337, September 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=2&spage=313.

Guo:2010:EGFb

[132] Li Guo and William Y. Sit. Enumeration and generating functions of differential Rota–Baxter words. Mathematics in Computer Science, 4(2– 3):339–358, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=2&spage=339.

Andres:2010:CMT

[133] Daniel Andres, Michael Brickenstein, Viktor Levandovskyy, Jorge Martín-Morales, and Hans Schönemann. Constructive D-module theory with Singular. Mathematics in Computer Science, 4(2–3):359–383, September 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=2&spage=359.

Vegter:2010:F

[134] Gert Vegter and Chee K. Yap. Foreword. Mathematics in Computer Science, 4(4):385–387, December 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=4&spage=385.

Gu:2010:FCC

[135] David Xianfeng Gu, Feng Luo, and Shing-Tung Yau. Fundamentals of computational conformal geometry. Mathematics in Computer Science, 4(4): 389–429, December 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=4&issue=4&spage=389.

Boissonnat:2010:TSS

[136] Jean-Daniel Boissonnat and Arijit Ghosh. Triangulating smooth submanifolds with light scaffolding. Mathematics in Computer Science, 4(4):431–461, December 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=4&issue=4&spage=
                                        431.

Yang:2010:EBA

[137] Huaiping Yang, Bert Jüttler, and Laureano Gonzalez-Vega. An evolutionbased approach for approximate parameterization of implicitly defined curves by polynomial parametric spline curves. Mathematics in Computer Science, 4 (4):463–479, December 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=4&spage=463.

Sagraloff:2010:GAI

[138] Michael Sagraloff. A general approach to isolating roots of a bitstream polynomial. Mathematics in Computer Science, 4(4):481–506, December 2010. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=4&spage=481.

Batra:2010:GCI

[139] Prashant Batra. Globally convergent, iterative path-following for algebraic equations. Mathematics in Computer Science, 4(4):507–537, December 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=4&issue=4&spage=507.

Morig:2010:DIG

[140] Marc Mörig, Ivo Rössling, and Stefan Schirra. On design and implementation of a generic number type for real algebraic number computations based on expression dags. Mathematics in Computer Science, 4(4):539–556, December 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=4&issue=4&spage=
539.

Miller:2011:F

[141] Mirka Miller, Bharati Rajan, and Joe Ryan. Foreword. Mathematics in Computer Science, 5(1):1–2, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=5&issue=1&spage=1.

Acharya:2011:SHP

[142] B. D. Acharya. Supra-hereditary properties of hypergraphs. Mathematics in Computer Science, 5(1):3–6, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=5&issue=1&spage=
                                        3.

Bommanahal:2011:PGC

[143] Basavanagoud Bommanahal and Keerthi G. Mirajkar. On Plick graphs with coarseness number one. Mathematics in Computer Science, 5(1):7–10, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=5&issue=1&spage=7.

Brankovic:2011:GLS

[144] Ljiljana Brankovic and Ian M. Wanless. Graceful labelling: State of the art, applications and future directions. Mathematics in Computer Science, 5(1):11–20, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=5&issue=1&spage=
                                        11.

Cada:2011:NDC

[145] Roman Cada, Evelyne Flandrin, and Haiyan Kang. A note on degree conditions for traceability in locally clawfree graphs. Mathematics in Computer Science, 5(1):21–25, March 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=1&spage=21.

Chaudhuri:2011:ISS

[146] Pranay Chaudhuri and Hussein Thompson. Improved self-stabilizing algorithms for L(2, 1)-labeling tree networks. Mathematics in Computer Science, 5(1):27–39, March 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=1&spage=27.

Ichishima:2011:POR

[147] Rikio Ichishima and Akito Oshima. On partitional and other related graphs. Mathematics in Computer Science, 5 (1):41–50, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=1&spage=41.

Jesintha:2011:AAF

[148] J. Jeba Jesintha and G. Sethuraman.

All arbitrarily fixed generalized banana trees are graceful. Mathematics in Computer Science, 5(1):51–62, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=1&spage=51.

Lopez:2011:SEM

[149] S. C. López, F. A. Muntaner-Batle, and M. Rius-Font. Super edge-magic models. Mathematics in Computer Science, 5(1):63–68, March 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=5&issue=1&spage=63.

ONeal:2011:ICO

[150] Allen O’Neal and Peter J. Slater. An introduction to closed/open neighborhood sums: Minimax, maximin, and spread. Mathematics in Computer Science, 5(1):69–80, March 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=5&issue=1&spage=69.

Phanalasy:2011:CAL

[151] Oudone Phanalasy, Mirka Miller, Costas S. Iliopoulos, Solon P. Pissis, and Elaheh Vaezpour. Construction of antimagic labeling for the cartesian product of regular graphs. Mathematics in Computer Science, 5(1):81–87, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=5&issue=1&spage=
                                        81.

Rajan:2011:CRH

[152] B. Rajan, K. T. Sonia, and M. Chris Monica. Conditional resolvability of honeycomb and hexagonal networks. Mathematics in Computer Science, 5(1):89–99, March 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=1&spage=89.

Rajasingh:2011:LWF

[153] Indra Rajasingh and Micheal Arockiaraj. Linear wirelength of folded hypercubes. Mathematics in Computer Science, 5(1):101–111, March 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=1&spage=101.

Sugeng:2011:CVM

[154] K. A. Sugeng and J. Ryan. Clique vertex magic cover of a graph. Mathematics in Computer Science, 5(1):113–118, March 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=5&issue=1&spage=
113.

Nishat:2011:VDP

[155] Rahnuma Islam Nishat, Debajyoti Mondal, and Md. Saidur Rahman. Visibility drawings of plane 3-trees with minimum area. Mathematics in Computer Science, 5(1):119–132, March 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=1&spage=119.

Manin:2011:ECC

[156] Yuri I. Manin and Matilde Marcolli.

Error-correcting codes and phase transitions. Mathematics in Computer Science, 5(2):133–170, June 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=2&spage=133.

Klein:2011:HCN

[157] Shmuel T. Klein and Dana Shapira. Huffman coding with non-sorted frequencies. Mathematics in Computer Science, 5(2):171–178, June 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=2&spage=171.

Faugere:2011:PER

[158] Jean-Charles Faugère and Ye Liang. Pivoting in extended rings for computing approximate Gröbner bases. Mathematics in Computer Science, 5 (2):179–194, June 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=2&spage=179.

Zhu:2011:SSF

[159] Mingfu Zhu, Guangran Jiang, and Shuhong Gao. Solving the 100 Swiss francs problem. Mathematics in Computer Science, 5(2):195–207, June 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=2&spage=195.

Gandy:2011:OPR

[160] Silvia Gandy, Masaaki Kanno, Hirokazu Anai, and Kazuhiro Yokoyama. Optimizing a particular real root of a polynomial by a special cylindrical algebraic decomposition. Mathematics in Computer Science, 5(2):209–221, June 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=2&spage=209.

Juttler:2011:QCS

[161] Bert Jüttler and Brian Moore. A quadratic clipping step with superquadratic convergence for bivariate polynomial systems. Mathematics in Computer Science, 5(2):223–235, June 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=5&issue=2&spage=
223.

Khattri:2011:OEO

[162] Sanjay Kumar Khattri. Optimal eighth order iterative methods. Mathematics in Computer Science, 5(2):237–243, June 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=5&issue=2&spage=
237.

Anai:2011:F

[163] Hirokazu Anai. Foreword. Mathematics in Computer Science, 5(3): 245–246, September 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=3&spage=245.

Li:2011:SAD

[164] Xiaoliang Li, Chenqi Mou, Wei Niu, and Dongming Wang. Stability analysis for discrete biological models using algebraic methods. Mathematics in Computer Science, 5(3):247–262, September 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=5&issue=3&spage=
247.

Mabrouki:2011:EBR

[165] Mbarka Mabrouki, Marc Aiguier, Jean-Paul Comet, Pascale Le Gall, and Adrien Richard. Embedding of biological regulatory networks and property preservation. Mathematics in Computer Science, 5(3):263–288, September 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=3&spage=263.

Boulier:2011:MRC

[166] François Boulier, Marc Lefranc, François Lemaire, and Pierre-Emmanuel Morant. Model reduction of chemical reaction systems using elimination. Mathematics in Computer Science, 5(3): 289–301, September 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=3&spage=289.

Ichihara:2011:SQA

[167] Hiroyuki Ichihara and Hirokazu Anai.

An SOS-QE approach to nonlinear gain analysis for polynomial dynamical systems. Mathematics in Computer Science, 5(3):303–314, September 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=3&spage=303.

Iwane:2011:SNA

[168] Hidenao Iwane, Hitoshi Yanami, and Hirokazu Anai. A symbolic-numeric approach to multi-objective optimization in manufacturing design. Mathematics in Computer Science, 5(3): 315–334, September 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=3&spage=315.

DosReis:2011:PCE

[169] Gabriel Dos Reis and Bjarne Stroustrup. A principled, complete, and efficient representation of C++. Mathematics in Computer Science, 5(3): 335–356, September 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=3&spage=335.

Fainekos:2011:F

[170] Georgios Fainekos, Eric Goubault, Sylvie Putot, and Stefan Ratschan. Foreword. Mathematics in Computer Science, 5(4):357–358, December 2011.

CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=4&spage=357.

Graillat:2011:SAM

[171] Stef Graillat, Fabienne Jézéquel, Shiyue Wang, and Yuxiang Zhu. Stochastic arithmetic in multiprecision. Mathematics in Computer Science, 5(4): 359–375, December 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=4&spage=359.

Boldo:2011:FVN

[172] Sylvie Boldo and Claude Marché. Formal verification of numerical programs: From C annotated programs to mechanical proofs. Mathematics in Computer Science, 5(4):377–393, December 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=4&spage=377.

Siegel:2011:TTA

[173] Stephen F. Siegel and Timothy K. Zirkel. TASS: The Toolkit for Accurate Scientific Software. Mathematics in Computer Science, 5(4): 395–426, December 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//vsl.cis.udel.edu/tass;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=4&spage=395.

Siegel:2011:FFE

[174] Stephen F. Siegel and Timothy K.

Zirkel. FEVS: a Functional Equivalence Verification Suite for high-performance scientific computing. Mathematics in Computer Science, 5(4):427–435, December 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1661-8270&volume=5&issue=4&spage=
427.

Collins:2011:VRF

[175] Pieter Collins, Milad Niqui, and Nathalie Revol. A validated real function calculus. Mathematics in Computer Science, 5(4):437–467, December 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=4&spage=437.

Damm:2011:PPV

[176] Werner Damm, Carsten Ihlemann, and Viorica Sofronie-Stokkermans. PTIME parametric verification of safety properties for reasonable linear hybrid automata. Mathematics in Computer Science, 5(4):469–497, December 2011. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=5&issue=4&spage=469.

Hellmuth:2012:SHP

[177] Marc Hellmuth, Lydia Ostermeier, and Peter F. Stadler. A survey on hypergraph products. Mathematics in Computer Science, 6(1):1–32, March 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=1&spage=1.

Rana:2012:CCP

[178] Akul Rana, Anita Pal, and Madhumangal Pal. The conditional covering problem on unweighted interval graphs with nonuniform coverage radius. Mathematics in Computer Science, 6(1):33–41, March 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=6&issue=1&spage=
                                        33.

Lundqvist:2012:MMI

[179] Samuel Lundqvist. Multiplication matrices and ideals of projective dimension zero. Mathematics in Computer Science, 6(1):43–59, March 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=1&spage=43.

Horn:2012:MFH

[180] Peter Horn, Wolfram Koepf, and Torsten Sprenger. m-fold hypergeometric solutions of linear recurrence equations revisited. Mathematics in Computer Science, 6(1):61–77, March 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=1&spage=61.

Nahay:2012:TOI

[181] John Michael Nahay. The n th order implicit differentiation formula for two variables with an application to computing all roots of a transcendental function. Mathematics in Computer Science, 6(1):79–105, March 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=1&spage=79.

Kotsireas:2012:F

[182] Ilias S. Kotsireas, Irene Márquez-Corbella, and Edgar Martínez-Moro. Foreword. Mathematics in Computer Science, 6(2):107–108, June 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=2&spage=107.

Jurrius:2012:RBM

[183] Relinde Jurrius. Relations between Möbius and coboundary polynomials. Mathematics in Computer Science, 6(2):109–120, June 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=2&spage=109.

Jurrius:2012:TFI

[184] Relinde Jurrius and Ruud Pellikaan.

Truncation formulas for invariant polynomials of matroids and geometric lattices. Mathematics in Computer Science, 6(2):121–133, June 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=2&spage=121.

Gordon:2012:BGD

[185] Gary Gordon. On Brylawski’s generalized duality. Mathematics in Computer Science, 6(2):135–146, June 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=2&spage=135.

Marquez-Corbella:2012:DMC

[186] Irene Márquez-Corbella and Edgar Martínez-Moro. Decomposition of modular codes for computing test sets and Graver basis. Mathematics in Computer Science, 6(2):147–165, June 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=2&spage=147.

Romanczuk:2012:FGL

[187] Urszula Romańczuk and Vasyl Ustimenko. On families of graphs of large cycle indicator, matrices of large order and key exchange protocols with nonlinear polynomial maps of small degree. Mathematics in Computer Science, 6(2):167–180, June 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=2&spage=167.

Klisowski:2012:CCP

[188] Michal Klisowski and Vasyl Ustimenko. On the comparison of cryptographical properties of two different families of graphs with large cycle indicator. Mathematics in Computer Science, 6(2):181–198, June 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=2&spage=181.

Marcolli:2012:CFN

[189] Matilde Marcolli and Christopher Perez.

Codes as fractals and noncommutative spaces. Mathematics in Computer Science, 6(3):199–215, September 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=3&spage=199.

Orange:2012:EAS

[190] Sébastien Orange, Guénaël Renault, and Kazuhiro Yokoyama. Efficient arithmetic in successive algebraic extension fields using symmetries. Mathematics in Computer Science, 6(3): 217–233, September 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=3&spage=217.

Kohlhase:2012:SOM

[191] Michael Kohlhase and Florian Rabe.

Semantics of OpenMath and MathML 3. Mathematics in Computer Science, 6(3):235–260, September 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=3&spage=235.

Aurenhammer:2012:CCH

[192] Franz Aurenhammer and Bert Jüttler. On computing the convex hull of (piecewise) curved objects. Mathematics in Computer Science, 6(3): 261–266, September 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=3&spage=261.

Li:2012:FSF

[193] Hongbo Li, Rida T. Farouki, and Dingkang Wang. Foreword to the special focus on mathematics and algorithms for CAM and CNC. Mathematics in Computer Science, 6(3):267–268, September 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1661-8270&volume=6&issue=3&spage=
                                        267.

Gasparetto:2012:TPR

[194] Alessandro Gasparetto, Paolo Boscariol, Albano Lanzutti, and Renato Vidoni. Trajectory planning in robotics. Mathematics in Computer Science, 6 (3):269–279, September 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=3&spage=269.

Li:2012:PSS

[195] Hongbo Li, Shoubin Yao, Ge Li, Yuanjie Liu, and Lixian Zhang. Power series solution for isoscallop tool path generation on free-form surface with ballend cutter. Mathematics in Computer Science, 6(3):281–296, September 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=3&spage=281.

Radzevich:2012:DKB

[196] Stephen P. Radzevich. The DG/ K-based approach for synthesizing of CAM system for sculptured surface machining on multi-axis NC machine. Mathematics in Computer Science, 6 (3):297–313, September 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=3&spage=297.

Wang:2012:CBR

[197] Yongqing Wang, Haibo Liu, and Sennan Yu. Curvature-based real-time NURBS surface interpolator with lookahead ACC/DEC control. Mathematics in Computer Science, 6(3): 315–326, September 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&
volume=6&issue=3&spage=315.

Li:2012:DIG

[198] Hongbo Li, Xiaoshan Gao, Lixian Zhang, and Ruiyong Sun. Discrete interpolation of G01 codes in 2D machining under bounded accelerations. Mathematics in Computer Science, 6 (3):327–344, September 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1661-8270&
                                        volume=6&issue=3&spage=327.

Ratschan:2012:F

[199] Stefan Ratschan and Thomas Sturm. Foreword. Mathematics in Computer Science, 6(4):345, December 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11786-012-0129-2;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s11786-012-0129-2.pdf.

Jaulin:2012:CIA

[200] Luc Jaulin. Combining interval analysis with flatness theory for state estimation of sailboat robots. Mathematics in Computer Science, 6(4):347–359, December 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-012-0130-9.

Huang:2012:RGS

[201] Yanli Huang, L. X. Châu Ngô, and Franz Winkler. Rational general solutions of trivariate rational differential systems. Mathematics in Computer Science, 6(4):361–374, December 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-012-0131-8.

Anashin:2012:NAT

[202] Vladimir Anashin. The non-Archimedean theory of discrete systems. Mathematics in Computer Science, 6(4):375– 393, December 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-012-
                                        0132-7.

Liu:2012:ADR

[203] Jiang Liu, Naijun Zhan, and Hengjun Zhao. Automatically discovering relaxed Lyapunov functions for polynomial dynamical systems. Mathematics in Computer Science, 6(4):395–408, December 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-012-0133-6.

Fietzke:2012:SDP

[204] Arnaud Fietzke and Christoph Weidenbach. Superposition as a decision procedure for timed automata. Mathematics in Computer Science, 6 (4):409–425, December 2012. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-012-0134-5.

Kruglov:2012:SDF

[205] Evgeny Kruglov and Christoph Weidenbach. Superposition decides the first-order logic fragment over ground theories. Mathematics in Computer Science, 6(4):427–456, December 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-012-0135-4.

Maza:2012:SPP

[206] Marc Moreno Maza, Bican Xia, and Rong Xiao. On solving parametric polynomial systems. Mathematics in Computer Science, 6(4):457–473, December 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-012-0136-3.

Barbosa:2012:CGQ

[207] L. S. Barbosa and Sun Meng. A calculus for generic, QoS-aware component composition. Mathematics in Computer Science, 6(4):475–497, December 2012. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-012-0137-2.

Naimpally:2013:F

[208] Som Naimpally, Jim Peters, and Marcin Wolski. Foreword. Mathematics in Computer Science, 7(1):1–2, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11786-013-0147-8;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s11786-013-0147-8.pdf.

Peters:2013:NSI

[209] James F. Peters. Near sets: An introduction. Mathematics in Computer Science, 7(1):3–9, March 2013. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s11786-013-0149-6;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s11786-013-0149-6.pdf.

Ramanna:2013:FGA

[210] Sheela Ramanna and Doungrat Chitcharoen. Flow graphs: Analysis with near sets. Mathematics in Computer Science, 7 (1):11–29, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-013-
                                        0144-y.

DiConcilio:2013:PFG

[211] A. Di Concilio. Point-free geometries: Proximities and quasi-metrics. Mathematics in Computer Science, 7 (1):31–42, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-013-
0140-2.

Fashandi:2013:NCU

[212] Homa Fashandi. Nearness of covering uniformities: Theory and application in image analysis. Mathematics in Computer Science, 7(1):43–50, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0142-0.

Henry:2013:MFN

[213] Christopher J. Henry. Metric free nearness measure using descriptionbased neighbourhoods. Mathematics in Computer Science, 7(1):51–69, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0141-1.

Henry:2013:SBP

[214] Christopher J. Henry and Sheela Ramanna. Signature-based perceptual nearness: Application of near sets to image retrieval. Mathematics in Computer Science, 7(1):71–85, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0145-x.

Peters:2013:LNS

[215] James F. Peters. Local near sets: Pattern discovery in proximity spaces. Mathematics in Computer Science, 7 (1):87–106, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-013-
0143-z.

Tiwari:2013:UCV

[216] Surabhi Tiwari. Ultrafilter completeness in ε-approach nearness spaces. Mathematics in Computer Science, 7(1): 107–111, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-013-
0148-7.

Wang:2013:GKM

[217] Lidong Wang, Xiaodong Liu, and Yashuang Mu. The global k-means clustering analysis based on multigranulations nearness neighborhood. Mathematics in Computer Science, 7(1): 113–124, March 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-013-
0150-0.

Wolski:2013:TFN

[218] Marcin Wolski. Toward foundations of near sets: (pre-)sheaf theoretic approach. Mathematics in Computer Science, 7(1):125–136, March 2013. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//link.springer.com/article/10.
1007/s11786-013-0146-9;   http://
link.springer.com/content/pdf/10.
1007/s11786-013-0146-9.pdf.

Conley:2013:IAB

[219] Ehud S. Conley and Shmuel T.

Klein. Improved alignment-based algorithm for multilingual text compression. Mathematics in Computer Science, 7(2):137–153, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-012-0138-1.

Cinque:2013:BIC

[220] Luigi Cinque, Sergio De Agostino, and Luca Lombardi. Binary image compression via monochromatic pattern substitution: Sequential and Parallel implementations. Mathematics in Computer Science, 7(2):155–166, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0153-x.

Perez-Diaz:2013:BFB

[221] Sonia Pérez-Díaz and J. Rafael Sendra.

Behavior of the fiber and the base points of parametrizations under projections. Mathematics in Computer Science, 7(2):167–184, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0139-8.

Zhao:2013:DVC

[222] Jun Zhao and Elizabeth Mansfield.

Discrete variational calculus for Bspline curves. Mathematics in Computer Science, 7(2):185–199, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0155-8.

Rosenkranz:2013:NAO

[223] M. Rosenkranz and A. Korporal. A noncommutative algebraic operational calculus for boundary problems. Mathematics in Computer Science, 7(2):201–227, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0154-9.

Li:2013:TPM

[224] Hongbo Li and Lixian Zhang. Two proofs on Max–Min–Max principle of jerk control in time-optimal rectilinear motion. Mathematics in Computer Science, 7(2):229–236, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-012-0128-3.

Siddiqui:2013:FAL

[225] Muhammad Kamran Siddiqui, Muhammad Numan, and Muhammad Awais Umar. Face antimagic labeling of Jahangir graph. Mathematics in Computer Science, 7(2):237–243, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0151-z.

Baca:2013:SFA

[226] Martin Baca, Muhammad Numan, and Muhammad Kamran Siddiqui. Super face antimagic labelings of union of antiprisms. Mathematics in Computer Science, 7(2):245–253, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0152-y.

Hellmuth:2013:PSP

[227] Marc Hellmuth, Wilfried Imrich, and Tomas Kupka. Partial star products: A local covering approach for the recognition of approximate Cartesian product graphs. Mathematics in Computer Science, 7(3):255–273, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0156-7.

Fagerberg:2013:CRC

[228] Rolf Fagerberg, Christoph Flamm, Daniel Merkle, Philipp Peters, and Peter F. Stadler. On the complexity of reconstructing chemical reaction networks. Mathematics in Computer Science, 7(3):275–292, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0160-y.

Elden:2013:CSC

[229] Lars Eldén, Magnus Merkel, Lars Ahrenberg, and Martin Fagerlund. Computing semantic clusters by semantic Mirroring and spectral graph partitioning. Mathematics in Computer Science, 7(3):293–313, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0159-4.

Ali:2013:MLT

[230] Kashif Ali, Muhammad Hussain, Ali Ahmad, and Mirka Miller. Magic labelings of type (a, b, c) of families of wheels. Mathematics in Computer Science, 7(3):315–319, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0162-9.

Semaev:2013:IAG

[231] Igor Semaev. Improved agreeing– gluing algorithm. Mathematics in Computer Science, 7(3):321–339, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0163-8.

Saeid:2013:STF

[232] Arsham Borumand Saeid, Akbar Rezaei, and Rajab Ali Borzooei. Some types of filters in BE-algebras. Mathematics in Computer Science, 7(3):341– 352, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-013-
                                        0157-6.

Akram:2013:IFS

[233] Muhammad Akram, Bijan Davvaz, and Feng Feng. Intuitionistic fuzzy soft K-algebras. Mathematics in Computer Science, 7(3):353–365, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0158-5.

Akram:2013:IFL

[234] Muhammad Akram, Saadia Shahzad, Arif Butt, and Abdul Khaliq. Intuitionistic fuzzy logic control for heater fans. Mathematics in Computer Science, 7(3):367–378, September 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0161-x.

Chen:2013:FSF

[235] Xiaoyu Chen, Dongming Wang, and Xiangliang Zhang. Foreword to the special focus on mathematics, data and knowledge. Mathematics in Computer Science, 7(4):379–386, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-013-0169-2.pdf.

Abbass:2013:EAI

[236] Mostafa M. Abbass and Hazem M.

Bahig. An efficient algorithm to identify DNA motifs. Mathematics in Computer Science, 7(4):387–399, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0165-6.

Cruz:2013:SNC

[237] Cristian Cruz, William Lima Leão, and David Rohde. The sensitivity of the number of clusters in a Gaussian mixture model to prior distributions. Mathematics in Computer Science, 7(4):401–420, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0168-3.

Aladova:2013:IMK

[238] Elena Aladova, Eugene Plotkin, and Tatjana Plotkin. Isotypeness of models and knowledge bases equivalence. Mathematics in Computer Science, 7(4):421–438, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-013-0166-5.

Chen:2013:FSG

[239] Xiaoyu Chen and Dongming Wang. Formalization and specification of geometric knowledge objects. Mathematics in Computer Science, 7(4):439– 454, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-013-
                                        0167-4.

Zima:2013:AIS

[240] Eugene V. Zima. Accelerating indefinite summation: Simple classes of summands. Mathematics in Computer Science, 7(4):455–472, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0170-9.

Roanes-Lozano:2013:GRG

[241] Eugenio Roanes-Lozano. The geometry of railway geometric overthrow revisited using computer algebra methods. Mathematics in Computer Science, 7(4):473–485, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-013-0164-7.

Baca:2013:TEI

[242] Martin Baca, Marcela Lascsáková, and Muhammad Kamran Siddiqui. Total edge irregularity strength of toroidal fullerene. Mathematics in Computer Science, 7(4):487–492, December 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0172-2.

Kerber:2014:F

[243] Manfred Kerber, Christoph Lange, and Colin Rowat. Foreword. Mathematics in Computer Science, 8(1): 1–4, March 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-014-0178-9.pdf.

Chatterjee:2014:ARS

[244] Siddharth Chatterjee and Arunava Sen.

Automated reasoning in social choice theory: Some remarks. Mathematics in Computer Science, 8(1):5–10, March 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0177-x.

James:2014:EFM

[245] Phillip James and Markus Roggenbach. Encapsulating formal methods within domain specific languages: A solution for verifying railway scheme plans. Mathematics in Computer Science, 8(1):11–38, March 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0174-0.

Khan-Afshar:2014:FAO

[246] Sanaz Khan-Afshar, Umair Siddique, Mohamed Yousri Mahmoud, Vincent Aravantinos, Ons Seddiki, Osman Hasan, and Sofiène Tahar. Formal analysis of optical systems. Mathematics in Computer Science, 8(1):39–70, March 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0175-z.

Mitsch:2014:CVD

[247] Stefan Mitsch, Grant Olney Passmore, and André Platzer. Collaborative verification-driven engineering of hybrid systems. Mathematics in Computer Science, 8(1):71–97, March 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0176-y.

Heras:2014:RPP

[248] Jónathan Heras and Ekaterina Komendantskaya. Recycling proof patterns in Coq: Case studies. Mathematics in Computer Science, 8(1):99–116, March 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0173-1.

DiRocco:2014:F

[249] Sandra Di Rocco and Josef Schicho. Foreword. Mathematics in Computer Science, 8(2):117–118, June 2014. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-014-0184-y.pdf.

Ren:2014:TCM

[250] Qingchun Ren, Steven V. Sam, and Bernd Sturmfels. Tropicalization of classical moduli spaces. Mathematics in Computer Science, 8(2):119–145, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0185-x.

Bremner:2014:HPF

[251] Murray Bremner, Jiaxiong Hu, and Luke Oeding. The 3 × 3 × 3 hyperdeterminant as a polynomial in the fundamental invariants for SL_{3}(C) × SL_{3}(C) × SL_{3}(C). Mathematics in Computer Science, 8(2):147–156, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0186-9.

Mourrain:2014:BDT

[252] Bernard Mourrain and Nelly Villamizar. Bounds on the dimension of trivariate spline spaces: A homological approach. Mathematics in Computer Science, 8(2):157–174, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0187-8.

Buse:2014:DSH

[253] Laurent Busé and Jean-Pierre Jouanolou.

On the discriminant scheme of homogeneous polynomials. Mathematics in Computer Science, 8(2):175–234, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0188-7.

Hauenstein:2014:NPW

[254] Jonathan D. Hauenstein and Frank Sottile. Newton polytopes and witness sets. Mathematics in Computer Science, 8(2):235–251, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0189-6.

Bates:2014:UMA

[255] Daniel J. Bates and Matthew Niemerg.

Using monodromy to avoid high precision in homotopy continuation. Mathematics in Computer Science, 8(2): 253–262, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-014-
0190-0.

Wilson:2014:CAS

[256] D. J. Wilson, R. J. Bradford, J. H. Davenport, and M. England. Cylindrical algebraic sub-decompositions. Mathematics in Computer Science, 8 (2):263–288, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-014-0191-z.pdf.

ElKahoui:2014:ACA

[257] M’hammed El Kahoui and Zakari Yaou Moussa. An algorithm to compute the adjoint ideal of an affine plane algebraic curve. Mathematics in Computer Science, 8(2):289–298, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0193-x.

Bastl:2014:PRR

[258] Bohumír Bastl, Bert Jüttler, Miroslav Lávicka, Tino Schulz, and Zbynek Sír. On the parameterization of rational ringed surfaces and rational canal surfaces. Mathematics in Computer Science, 8(2):299–319, June 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0192-y.

Ramdani:2014:F

[259] Nacim Ramdani and Luc Jaulin. Foreword. Mathematics in Computer Science, 8(3–4):321, September 2014. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-014-0211-z.pdf.

Rohn:2014:VLD

[260] Jiri Rohn. Verification of linear (in)dependence in finite precision arithmetic. Mathematics in Computer Science, 8(3–4):323–328, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0196-7.

Hladik:2014:ESC

[261] Milan Hladík and Stefan Ratschan. Efficient solution of a class of quantified constraints with quantifier prefix exists–forall. Mathematics in Computer Science, 8(3–4):329–340, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0195-8.

Just:2014:SSB

[262] Elke Just. Subdivision strategies for boxes in branch-and-bound nonlinear solvers and verification. Mathematics in Computer Science, 8(3–4):341– 355, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-014-
0207-8.

Patil:2014:IBG

[263] Bhagyesh V. Patil and P. S. V. Nataraj. An improved Bernstein global optimization algorithm for MINLP problems with application in process industry. Mathematics in Computer Science, 8(3–4):357–377, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0198-5.

Aubry:2014:KCI

[264] Clément Aubry, Rozenn Desmare, and Luc Jaulin. Kernel characterization of an interval function. Mathematics in Computer Science, 8(3–4):379– 390, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-014-
0206-9.

Ravanbod:2014:DSM

[265] Laleh Ravanbod, Nathalie Verdière, and Carine Jauberthie. Determination of set–membership identifiability sets. Mathematics in Computer Science, 8(3–4):391–406, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0201-1.

Maiga:2014:CVZ

[266] Moussa Maïga, Nacim Ramdani, Louise Travé-Massuyès, and Christophe Combastel. A CSP versus a zonotopebased method for solving guard set intersection in nonlinear hybrid reachability. Mathematics in Computer Science, 8(3–4):407–423, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0204-y.

Rego:2014:DIO

[267] Francisco Rego, Elwin de Weerdt, Eddy van Oort, Erik-Jan van Kampen, Qiping Chu, and António M. Pascoal. Determination of inner and outer bounds of reachable sets through subpavings. Mathematics in Computer Science, 8(3–4):425–442, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0199-4.

Kumkov:2014:LSV

[268] Sergey S. Kumkov, Stéphane Le Ménec, and Valerii S. Patsko. Level sets of the value function in differential games with two pursuers and one evader. interval analysis interpretation. Mathematics in Computer Science, 8(3–4):443–454, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0203-z.

Chhade:2014:NPD

[269] Hiba Haj Chhadé, Amadou Gning, Fahed Abdallah, Imad Mougharbel, and Simon Julier. Non parametric distributed inference in sensor networks using box particles messages. Mathematics in Computer Science, 8(3–4): 455–478, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-014-0200-2.pdf.

Sandretto:2014:IMM

[270] Julien Alexandre Dit Sandretto, Gilles Trombettoni, and David Daney. Interval methods for model qualification: Methodology and advanced application. Mathematics in Computer Science, 8(3–4):479–493, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0210-0.

Seddik:2014:PBL

[271] Mohamed Saad Ibn Seddik, Luc Jaulin, and Jonathan Grimsdale. Phase based localization for underwater vehicles using interval analysis. Mathematics in Computer Science, 8(3–4):495– 502, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-014-
                                        0197-6.

Bethencourt:2014:SNL

[272] Aymeric Bethencourt and Luc Jaulin. Solving non-linear constraint satisfaction problems involving time-dependent functions. Mathematics in Computer Science, 8(3–4):503–523, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0209-6.

Rauh:2014:IMR

[273] Andreas Rauh, Luise Senkel, Ekaterina Auer, and Harald Aschemann. Interval methods for real-time capable robust control of solid oxide fuel cell systems. Mathematics in Computer Science, 8(3–4):525–542, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0205-x.

Senkel:2014:SMT

[274] Luise Senkel, Andreas Rauh, and Harald Aschemann. Sliding mode techniques for robust trajectory tracking as well as state and parameter estimation. Mathematics in Computer Science, 8(3–4):543–561, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0208-7.

Drevelle:2014:VVI

[275] Vincent Drevelle and Jeremy Nicola.

VIBes: A visualizer for intervals and boxes. Mathematics in Computer Science, 8(3–4):563–572, September 2014. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0202-0.

Thery:2015:FSF

[276] Laurent Théry and Freek Wiedijk. Foreword to the special focus on formal proofs for mathematics and computer science. Mathematics in Computer Science, 9(1):1–3, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-014-0214-9.pdf.

Kaliszyk:2015:HHO

[277] Cezary Kaliszyk and Josef Urban.

HOL(y)Hammer: Online ATP service for HOL light. Mathematics in Computer Science, 9(1):5–22, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-014-0182-0.pdf.

Noschinski:2015:GLI

[278] Lars Noschinski. A graph library for Isabelle. Mathematics in Computer Science, 9(1):23–39, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0183-z.

Boldo:2015:CUF

[279] Sylvie Boldo, Catherine Lelay, and Guillaume Melquiond. Coquelicot: A user-friendly library of real analysis for Coq. Mathematics in Computer Science, 9(1):41–62, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0181-1.

Marcolli:2015:QCR

[280] Matilde Marcolli and John Napp. Quantum computation and real multiplication. Mathematics in Computer Science, 9(1):63–84, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-014-0179-8.

Talwar:2015:SSA

[281] Jyoti Talwar and Ranjan Kumar Mohanty. A single sweep AGE algorithm based on off-step discretization for the solution of viscous Burgers’ equation on a variable mesh. Mathematics in Computer Science, 9(1):85–103, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0194-9.

Daykin:2015:AGC

[282] Jacqueline W. Daykin, Costas S. Iliopoulos, Mirka Miller, and Oudone Phanalasy. Antimagicness of generalized corona and snowflake graphs. Mathematics in Computer Science, 9(1): 105–111, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-014-
0213-x.

Paul:2015:LPB

[283] Satyabrata Paul, Madhumangal Pal, and Anita Pal. L(2, 1)-labeling of permutation and bipartite permutation graphs. Mathematics in Computer Science, 9(1):113–123, March 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0180-2.

Miller:2015:F

[284] Mirka Miller, Bharati Rajan, and Indra Rajasingh. Foreword. Mathematics in Computer Science, 9 (2):125–126, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

//link.springer.com/article/10.
1007/s11786-015-0235-z;   http://
link.springer.com/content/pdf/10.
1007/s11786-015-0235-z.pdf.

Abawajy:2015:DCS

[285] J. Abawajy, A. V. Kelarev, M. Miller, and J. Ryan. Distances of centroid sets in a graph-based construction for information security applications. Mathematics in Computer Science, 9(2):127–137, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-015-0217-1.

Baca:2015:ALJ

[286] Martin Baca, Oudone Phanalasy, and Joe Ryan. Antimagic labelings of join graphs. Mathematics in Computer Science, 9(2):139–143, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-015-0218-0.

Conde:2015:NAM

[287] Josep Conde, Mirka Miller, Josep M. Miret, and Kumar Saurav. On the nonexistence of almost Moore digraphs of degree four and five. Mathematics in Computer Science, 9(2):145–149, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-015-0219-z.

Rajasingh:2015:TVI

[288] Indra Rajasingh and V. Annamma. Total vertex irregularity strength of 1-fault tolerant Hamiltonian graphs. Mathematics in Computer Science, 9(2): 151–160, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-015-
                                        0220-6.

Indriati:2015:TEI

[289] Diari Indriati, Widodo, and Indah E. Wijayanti. On total edge irregularity strength of generalized Web graphs and related graphs. Mathematics in Computer Science, 9(2):161–167, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0221-5.

Quadras:2015:DPC

[290] Jasintha Quadras and Sajiya Merlin Mahizl Albert. Domination parameters in coronene torus network. Mathematics in Computer Science, 9 (2):169–175, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-015-
0222-4.

Quadras:2015:EFH

[291] Jasintha Quadras and Sarah Surya Solomon. Embedding of the folded hypercubes into tori. Mathematics in Computer Science, 9(2):177–183, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0223-3.

Jesintha:2015:AUB

[292] J. Jeba Jesintha and K. Ezhilarasi Hilda. All uniform bow graphs are graceful. Mathematics in Computer Science, 9(2):185–191, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0224-2.

Punitha:2015:SKN

[293] M. Joice Punitha. Strong kernel number in certain oriented cycle extension of graphs. Mathematics in Computer Science, 9(2):193–199, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0225-1.

Manuel:2015:SMD

[294] Paul Manuel, Bharati Rajan, and Cyriac Grigorious. On the strong metric dimension of tetrahedral diamond lattice. Mathematics in Computer Science, 9(2):201–208, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0226-0.

Manuel:2015:BFC

[295] Paul Manuel and A. S. Shanthi. Berge–Fulkerson conjecture on certain snarks. Mathematics in Computer Science, 9(2):209–220, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0227-z.

Rajasingh:2015:LLD

[296] Indra Rajasingh, Micheal Arockiaraj, and Jasintha Quadras. Linear layout of directed grid graph. Mathematics in Computer Science, 9(2):221–227, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0228-y.

Ramdani:2015:TIS

[297] Rismawati Ramdani, A. N. M. Salman, and Hilda Assiyatun. Total irregularity strength of three families of graphs. Mathematics in Computer Science, 9(2):229–237, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0229-x.

Pushpam:2015:SRD

[298] P. Roushini Leely Pushpam and Chitra Suseendran. Secure restrained domination in graphs. Mathematics in Computer Science, 9(2):239–247, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0230-4.

Sethuraman:2015:CPT

[299] G. Sethuraman, A. Velankanni, and S. Anbarasu. Cycle partition of two-connected and two-edge connected graphs. Mathematics in Computer Science, 9(2):249–252, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0231-3.

Rajan:2015:ERC

[300] R. Sundara Rajan, N. Parthiban, and T. M. Rajalaxmi. Embedding of recursive circulants into certain necklace graphs. Mathematics in Computer Science, 9(2):253–263, June 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0232-2.

Fontaine:2015:FSF

[301] Pascal Fontaine, Thomas Sturm, and Uwe Waldmann. Foreword to the special focus on constraints and combinations. Mathematics in Computer Science, 9(3):265, October 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-015-0239-8.pdf.

Fukasaku:2015:QAA

[302] Ryoya Fukasaku, Shutaro Inoue, and Yosuke Sato. On QE algorithms over an algebraically closed field based on comprehensive Gröbner systems. Mathematics in Computer Science, 9(3):267–281, October 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0237-x.

Hagemann:2015:EGO

[303] Willem Hagemann. Efficient geometric operations on convex polyhedra, with an application to reachability analysis of hybrid systems. Mathematics in Computer Science, 9(3):283–325, October 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0238-9.

Suda:2015:VCE

[304] Martin Suda. Variable and clause elimination for LTL satisfiability checking. Mathematics in Computer Science, 9(3):327–344, October 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0240-2.

Wintraecken:2015:OTC

[305] M. H. M. J. Wintraecken and G. Vegter. On the optimal triangulation of convex hypersurfaces, whose vertices lie in ambient space. Mathematics in Computer Science, 9(3):345–353, October 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0216-7.

Panayotopoulos:2015:PMC

[306] A. Panayotopoulos and P. Vlamos. Partitioning the meandering curves. Mathematics in Computer Science, 9(3): 355–364, October 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-015-
0234-0.

Koepf:2015:AAF

[307] Wolfram Koepf and Etienne Nana Chiadjeu. Algorithmic approach for formal Fourier series. Mathematics in Computer Science, 9(3):365–389, October 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-014-0215-8.

Marcolli:2015:GGI

[308] Matilde Marcolli and Alexander Port.

Graph grammars, insertion Lie algebras, and quantum field theory. Mathematics in Computer Science, 9(4):391– 408, December 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-015-
0236-y.

Stadler:2015:CS

[309] Bärbel M. R. Stadler and Peter F.

Stadler. Connectivity spaces. Mathematics in Computer Science, 9(4):409– 436, December 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-015-
0241-1.

Berkemer:2015:SCR

[310] Sarah J. Berkemer, Ricardo R. C. Chaves, and Adrian Fritz. Spiders can be recognized by counting their legs. Mathematics in Computer Science, 9(4):437–441, December 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0233-1.

Muhshi:2015:MSR

[311] Hadi Muhshi and Edy Tri Baskoro. Matching-star Ramsey minimal graphs. Mathematics in Computer Science, 9 (4):443–452, December 2015. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0244-y.

Man:2015:CMA

[312] Yiu-Kwong Man. On computing the measurable amounts of the two jugs problem. Mathematics in Computer Science, 9(4):453–459, December 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-015-0242-0.

Kumar:2015:LBC

[313] Sandeep Ameet Kumar, Jito Vanualailai, and Bibhya Sharma. Lyapunovbased control for a swarm of planar nonholonomic vehicles. Mathematics in Computer Science, 9(4):461– 475, December 2015. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-015-
0243-z.

Sendra:2016:F

[314] J. Rafael Sendra, Dongming Wang, and Jing Yang. Foreword. Mathematics in Computer Science, 10(1): 1–3, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-016-0263-3.pdf.

Botana:2016:UUT

[315] Francisco Botana and Tomas Recio. On the unavoidable uncertainty of truth in dynamic geometry proving. Mathematics in Computer Science, 10(1):5–25, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0246-4.

Schreck:2016:CCT

[316] Pascal Schreck, Vesna Marinković, and Predrag Janicić. Constructibility classes for triangle location problems. Mathematics in Computer Science, 10(1):27–39, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0255-3.

Schreck:2016:ACC

[317] Pascal Schreck and Pascal Mathis.

Automatic constructibility checking of a corpus of geometric construction problems. Mathematics in Computer Science, 10(1):41–56, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0247-3.

Narboux:2016:TCV

[318] Julien Narboux and David Braun. Towards a certified version of the encyclopedia of triangle centers. Mathematics in Computer Science, 10(1):57–73, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-016-0254-4.

Shao:2016:CTP

[319] Changpeng Shao, Hongbo Li, and Lei Huang. Challenging theorem provers with mathematical Olympiad problems in solid geometry. Mathematics in Computer Science, 10(1):75–96, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-016-0256-2.

Dorst:2016:CCM

[320] Leo Dorst. The construction of 3D conformal motions. Mathematics in Computer Science, 10(1):97– 113, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-016-0250-8.pdf.

Bowers:2016:GUM

[321] John C. Bowers and Ileana Streinu. Geodesic universal molecules. Mathematics in Computer Science, 10(1): 115–141, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-016-
                                        0253-5.

Lichtblau:2016:FOP

[322] Daniel Lichtblau. First order perturbation and local stability of parametrized systems. Mathematics in Computer Science, 10(1):143–163, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0249-1.

Brunat:2016:CCR

[323] Josep M. Brunat and Antonio Montes.

Computing the canonical representation of constructible sets. Mathematics in Computer Science, 10(1): 165–178, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-016-
0248-2.

Ricca:2016:DCS

[324] Giorgio Ricca, Mauro C. Beltrametti, and Anna Maria Massone. Detecting curves of symmetry in images via Hough transform. Mathematics in Computer Science, 10(1):179–205, March 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0245-5.

Rajan:2016:FSF

[325] Bharati Rajan, Ibrahim Venkat, and K. G. Subramanian. Foreword to the special focus on graph theory and applications. Mathematics in Computer Science, 10(2):207–208, June 2016. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-016-0272-2.pdf.

Bera:2016:SPW

[326] Somnath Bera and Kalpana Mahalingam. Structural properties of word representable graphs. Mathematics in Computer Science, 10(2):209–222, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-016-0257-1.

Wong:2016:CBC

[327] Denis C. K. Wong. Constructions of binary codes based on bipartite graphs. Mathematics in Computer Science, 10(2):223–227, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-016-0258-0.

Nagar:2016:ECI

[328] Atulya K. Nagar and S. Sriram. On eccentric connectivity index of eccentric graph of regular dendrimer. Mathematics in Computer Science, 10(2): 229–237, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-016-
                                        0259-z.

Sivagami:2016:CCN

[329] P. Sivagami and Indra Rajasingh. T - coloring of certain networks. Mathematics in Computer Science, 10(2): 239–248, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s11786-016-
                                        0260-6.

Mahmood:2016:MCE

[330] Ali Abdulkareem Mahmood, Ali Maroosi, and Ravie Chandren Muniyandi. Membrane computing to enhance time efficiency of minimum dominating set. Mathematics in Computer Science, 10(2):249–261, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0261-5.

Sivaselvan:2016:IMQ

[331] K. Sivaselvan and C. Vijayalakshmi.

Implementation of Markovian queueing network model with multiple closed chains. Mathematics in Computer Science, 10(2):263–272, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0262-4.

Dochviri:2016:TSF

[332] Irakli Dochviri and James F. Peters. Topological sorting of finitely near sets. Mathematics in Computer Science, 10(2):273–277, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0273-1.

Gavina:2016:IAC

[333] Alexandra Gavina, José Matos, and Paulo Vasconcelos. Improving the accuracy of Chebyshev Tau method for nonlinear differential problems. Mathematics in Computer Science, 10(2): 279–289, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s11786-016-
                                        0265-1.

Conceicao:2016:ESS

[334] Ana C. Conceição and José C. Pereira. Exploring the spectra of some classes of singular integral operators with symbolic computation. Mathematics in Computer Science, 10(2):291–309, June 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-016-0264-2.

Loja:2016:FSF

[335] Amélia Loja, José Alberto Rodrigues, and Ana C. Conceição. Foreword to the special focus on advances in symbolic and numeric computation. Mathematics in Computer Science, 10(3): 311–312, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-016-0270-4.pdf.

Trindade:2016:TLM

[336] M. Trindade, J. Matos, and P. B. Vasconcelos. Towards a Lanczos’ τmethod toolkit for differential problems. Mathematics in Computer Science, 10(3):313–329, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s11786-016-0269-x.

Rachah:2016:DOC

[337] Amira Rachah and Delfim F. M. Torres. Dynamics and optimal control of Ebola transmission. Mathematics in Computer Science, 10(3):331–342, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0268-y.

Maia:2016:TDC

[338] Rúben Maia, Ricardo Branco, F. V.

Antunes, M. C. Oliveira, and Andrei Kotousov. Three-dimensional computational analysis of stress state transition in through-cracked plates. Mathematics in Computer Science, 10(3):343– 352, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-016-
0267-z.

Escobar:2016:CLA

[339] J. M. Escobar, J. Núñez, and P. Pérez-Fernández. On contractions of Lie algebras. Mathematics in Computer Science, 10(3):353–364, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0266-0.

Conceicao:2016:SCA

[340] Ana C. Conceição, Rui C. Marreiros, and José C. Pereira. Symbolic computation applied to the study of the kernel of a singular integral operator with non–Carleman shift and conjugation. Mathematics in Computer Science, 10(3):365–386, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0271-3.

Akram:2016:PFL

[341] Muhammad Akram and Arooj Adeel.

m⃗-polar fuzzy labeling graphs with application. Mathematics in Computer Science, 10(3):387–402, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0277-x.

Hammack:2016:ANA

[342] Richard H. Hammack, Marc Hellmuth, Lydia Ostermeier, and Peter F. Stadler. Associativity and nonassociativity of some hypergraph products. Mathematics in Computer Science, 10(3):403–408, September 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0276-y.

Guidi:2016:SRM

[343] Ferruccio Guidi and Claudio Sacerdoti Coen. A survey on retrieval of mathematical knowledge. Mathematics in Computer Science, 10(4):409– 427, December 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-016-
0274-0.

Raggi:2016:ACR

[344] Daniel Raggi, Alan Bundy, Gudmund Grov, and Alison Pease. Automating change of representation for proofs in discrete mathematics (extended version). Mathematics in Computer Science, 10(4):429–457, December 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-016-0275-z.pdf.

Manin:2016:SS

[345] Yuri I. Manin and Matilde Marcolli. Semantic spaces. Mathematics in Computer Science, 10(4):459–477, December 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-016-0278-9.

Wang:2016:CBR

[346] Dongming Wang. On the connection between ritt characteristic sets and Buchberger–Gröbner bases. Mathematics in Computer Science, 10(4):479– 492, December 2016. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-016-
0279-8.

Kotsireas:2017:MCS

[347] Ilias S. Kotsireas. Mathematics in computer science: After 10 years. Mathematics in Computer Science, 11(1):1–6, March 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-017-0312-6.pdf.

Chen:2017:AOR

[348] Yuqun Chen, Haibin Wu, and Honglian Xie. Automaticity of one–relator semigroups with length less than or equal to three. Mathematics in Computer Science, 11(1):7–33, March 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-017-0291-7.

Emiris:2017:AMS

[349] Ioannis Z. Emiris, Anna Karasoulou, and Charilaos Tzovas. Approximating multidimensional subset sum and Minkowski decomposition of polygons. Mathematics in Computer Science, 11(1):35–48, March 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-017-0297-1.

Galan-Garcia:2017:GAA

[350] José Luis Galán-García, Salvador Merino, Javier Martínez, and Miguel De Aguilera. Genetic and algebraic algorithms for classifying the items of a Likert questionnaire. Mathematics in Computer Science, 11(1):49–59, March 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-017-0289-1.

Hashemi:2017:GSC

[351] Amir Hashemi, Mahdi Dehghani Darmian, and Marzieh Barkhordar. Gröbner systems conversion. Mathematics in Computer Science, 11(1): 61–77, March 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11786-017-
0295-3.

Shu:2017:SSC

[352] Kevin Shu and Matilde Marcolli. Syntactic structures and code parameters. Mathematics in Computer Science, 11(1):79–90, March 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-017-0298-0.

Yang:2017:STR

[353] Fan Yang, Xiao-Xiao Li, Dun-Gang Li, and Lan Wang. The simplified Tikhonov regularization method for solving a Riesz–Feller space–fractional backward diffusion problem. Mathematics in Computer Science, 11(1):91–110, March 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL

http://link.springer.com/article/
10.1007/s11786-017-0292-6.

Iliopoulos:2017:F

[354] Costas S. Iliopoulos and Alessio Langiu.

Foreword. Mathematics in Computer Science, 11(2):111–112, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-017-0323-3.pdf.

BenNsira:2017:LSM

[355] Nadia Ben Nsira, Mourad Elloumi, and Thierry Lecroq. On-line string matching in highly similar DNA sequences. Mathematics in Computer Science, 11 (2):113–126, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kim:2017:RTO

[356] Jinil Kim, Amihood Amir, Joong Chae Na, Kunsoo Park, and Jeong Seop Sim. On representations of ternary order relations in numeric strings. Mathematics in Computer Science, 11(2):127–136, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Karkkainen:2017:ELE

[357] Juha Kärkkäinen and Dominik Kempa. Engineering a lightweight external memory suffix array construction algorithm. Mathematics in Computer Science, 11 (2):137–149, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Gagie:2017:CSS

[358] Travis Gagie, Giovanni Manzini, and Daniel Valenzuela. Compressed spaced suffix arrays. Mathematics in Computer Science, 11(2):151–157, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Tischler:2017:FAC

[359] German Tischler. Faster average case low memory semi-external construction of the Burrows–Wheeler transform. Mathematics in Computer Science, 11(2):159–176, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-017-0296-2.pdf.

Davoodi:2017:SRB

[360] Pooya Davoodi, Rajeev Raman, and Srinivasa Rao Satti. On succinct representations of binary trees. Mathematics in Computer Science, 11(2):177–189, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Gagie:2017:BGP

[361] Travis Gagie, Christopher Hoobin, and Simon J. Puglisi. Block graphs in practice. Mathematics in Computer Science, 11(2):191–196, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Vlachakis:2017:APA

[362] Dimitrios Vlachakis, Alexandros Armaos, and Sophia Kossida. Advanced protein alignments based on sequence, structure and hydropathy profiles; the paradigm of the viral polymerase enzyme. Mathematics in Computer Science, 11(2):197–208, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Daykin:2017:ISF

[363] Jacqueline W. Daykin and Bruce Watson. Indeterminate string factorizations and degenerate text transformations. Mathematics in Computer Science, 11(2):209–218, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-016-0285-x.pdf.

Hasan:2017:PSA

[364] Md. Mahbubul Hasan, A. S. M. Sohidull Islam, M. Sohel Rahman, and Ayon Sen. Palindromic subsequence automata and longest common palindromic subsequence. Mathematics in Computer Science, 11(2):219–232, June 2017. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Ferdous:2017:SMC

[365] S. M. Ferdous and M. Sohel Rahman.

Solving the minimum common string partition problem with the help of ants. Mathematics in Computer Science, 11 (2):233–249, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Dana-Picard:2017:F

[366] Thierry Dana-Picard, Wolfram Koepf, Ilias Kotsireas, Zoltán Kovács, Alexander Prokopenya, and Werner Seiler. Foreword. Mathematics in Computer Science, 11(3–4):251–252, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s11786-
017-0324-2.pdf.

Bavula:2017:QGW

[367] V. V. Bavula. Quiver generalized Weyl algebras, skew category algebras and diskew polynomial rings. Mathematics in Computer Science, 11(3– 4):253–268, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-017-0313-5.pdf.

Hussein:2017:GII

[368] Wuria Muhammad Ameen Hussein and Colin Christopher. A geometric investigation of the invariant algebraic curves in two dimensional Lotka–Volterra systems. Mathematics in Computer Science, 11(3–4):269–283, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Dana-Picard:2017:FIB

[369] Thierry Dana-Picard and David G. Zeitoun. A framework for an ICT–Based study of parametric integrals. Mathematics in Computer Science, 11(3–4): 285–296, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Davenport:2017:WDL

[370] James H. Davenport. What does “without loss of generality” mean, and how do we detect it. Mathematics in Computer Science, 11(3–4): 297–303, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-017-0316-2.pdf.

Durcheva:2017:ACT

[371] Mariana Durcheva and Elena Varbanova. Applications of CAS in the teaching and learning of discrete mathematics. Mathematics in Computer Science, 11(3–4):305–314, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Duzhin:2017:MAC

[372] Vasilii Duzhin and Nikolay Vasilyev.

Modeling of an asymptotically central Markov process on 3D Young graph. Mathematics in Computer Science, 11 (3–4):315–328, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Furst:2017:RRB

[373] Christoph Fürst and Alexander Levin.

Relative reduction and Buchberger’s algorithm in filtered free modules. Mathematics in Computer Science, 11(3– 4):329–339, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-017-0317-1.pdf.

Jeffrey:2017:BSI

[374] David J. Jeffrey. Branch structure and implementation of Lambert W. Mathematics in Computer Science, 11(3–4): 341–350, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kovacs:2017:RTA

[375] Zoltán Kovács. Real-time animated dynamic geometry in the classrooms by using fast Gröbner basis computations. Mathematics in Computer Science, 11 (3–4):351–361, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Wojas:2017:FSD

[376] Wlodzimierz Wojas and Jan Krupa. Familiarizing students with definition of Lebesgue integral: Examples of calculation directly from its definition using Mathematica. Mathematics in Computer Science, 11(3–4): 363–381, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-017-0321-5.pdf.

Minglibayev:2017:TBP

[377] M. Zh. Minglibayev, A. N. Prokopenya, G. M. Mayemerova, and Zh. U. Imanova. Three-body problem with variable masses that change anisotropically at different rates. Mathematics in Computer Science, 11(3–4):383–391, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Oldenburg:2017:TRB

[378] Reinhard Oldenburg. Transparent rule based CAS to support formalization of knowledge. Mathematics in Computer Science, 11(3–4):393–399, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Farrington:2017:SCR

[379] Eleanor Farrington and Emma Previato. Symbolic computation for Rankin– Cohen differential algebras: A case study. Mathematics in Computer Science, 11(3–4):401–415, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Prokopenya:2017:MSA

[380] Alexander N. Prokopenya. Motion of a swinging Atwood’s machine: Simulation and analysis with Mathematica. Mathematics in Computer Science, 11 (3–4):417–425, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-017-0301-9.pdf.

Quaresma:2017:TID

[381] Pedro Quaresma. Towards an intelligent and dynamic geometry book. Mathematics in Computer Science, 11(3–4): 427–437, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Roanes-Lozano:2017:BNA

[382] Eugenio Roanes-Lozano. A brief note on the approach to the conic sections of a right circular cone from dynamic geometry. Mathematics in Computer Science, 11(3–4):439–448, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Sarafian:2017:ACA

[383] Haiduke Sarafian. Application of computer algebra system and the meanvalue theory for evaluating electrostatic potential and its associated field for nontrivial configurations. Mathematics in Computer Science, 11(3–4):449–455, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Siluszyk:2017:CCC

[384] Agnieszka Siluszyk. On a class of central configurations in the planar 3n-body problem. Mathematics in Computer Science, 11(3–4):457–467, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-017-0309-1.pdf.

Chrapary:2017:DCS

[385] Hagen Chrapary, Wolfgang Dalitz, Winfried Neun, and Wolfram Sperber. Design, concepts, and state of the art of the swMATH service. Mathematics in Computer Science, 11(3–4):469–481, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Noro:2018:UMT

[386] Masayuki Noro and Kazuhiro Yokoyama. Usage of modular techniques for efficient computation of ideal operations. Mathematics in Computer Science, 12(1):1–32, March 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Port:2018:PTS

[387] Alexander Port, Iulia Gheorghita, Daniel Guth, John M. Clark, Crystal Liang, Shival Dasu, and Matilde Marcolli. Persistent topology of syntax. Mathematics in Computer Science, 12 (1):33–50, March 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kleine:2018:EDI

[388] Kristoffer Kleine and Dimitris E. Simos.

An efficient design and implementation of the in–parameter-order algorithm. Mathematics in Computer Science, 12 (1):51–67, March 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Wang:2018:UDQ

[389] Haohao Wang and Ron Goldman. Using dual quaternion to study translational surfaces. Mathematics in Computer Science, 12(1):69–75, March 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Lu:2018:ALL

[390] Yingyu Lu, Guanghua Dong, Wenhui Ma, and Ning Wang. Antimagic labeling of the lexicographic product graph K_{m,n}[P_{k}]. Mathematics in Computer Science, 12(1):77–90, March 2018. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

DiConcilio:2018:DPP

[391] A. Di Concilio, C. Guadagni, J. F.

Peters, and S. Ramanna. Descriptive proximities. properties and interplay between classical proximities and overlap. Mathematics in Computer Science, 12 (1):91–106, March 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Loja:2018:FSF

[392] Amélia Loja, Stéphane Louis Clain, Joaquim Infante Barbosa, and José Alberto Rodrigues. Foreword to the special focus on advances in symbolic and numeric computation II.

Mathematics in Computer Science, 12 (2):107–109, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-018-0341-9.pdf.

Allali:2018:AOC

[393] Karam Allali, Sanaa Harroudi, and Delfim F. M. Torres. Analysis and optimal control of an intracellular delayed HIV model with CTL immune response. Mathematics in Computer Science, 12 (2):111–127, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Mendes:2018:PCM

[394] I. R. Mendes and P. B. Vasconcelos. PageRank computation with MAAOR and lumping methods. Mathematics in Computer Science, 12(2):129–141, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Escobar:2018:NOP

[395] J. M. Escobar, J. Núñez, and P. Pérez-Fernández. A new one-parameter invariant function for algebras. Mathematics in Computer Science, 12(2):143–150, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Macedo:2018:SAG

[396] Ângela Macedo, Teresa A. Mesquita, and Zélia da Rocha. Symbolic approach to the general quadratic polynomial decomposition. Mathematics in Computer Science, 12(2):151–172, June 2018. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Ralha:2018:MPB

[397] Rui Ralha. Mixed precision bisection.

Mathematics in Computer Science, 12 (2):173–181, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Trindade:2018:DFC

[398] M. Trindade, J. Matos, and P. B. Vasconcelos. Dealing with functional coefficients within Tau method. Mathematics in Computer Science, 12(2):183–195, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Matos:2018:SDI

[399] J. C. Matos, J. M. A. Matos, and M. J. Rodrigues. Solving differential and integral equations with Tau method. Mathematics in Computer Science, 12 (2):197–205, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Retzlaff:2018:PLM

[400] Nancy Retzlaff and Peter F. Stadler.

Partially local multi-way alignments. Mathematics in Computer Science, 12 (2):207–234, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-018-0338-4.pdf.

Artinescu:2018:SCA

[401] Irina Maria Artinescu and Liviu Octavian Mafteiu-Scai. A scratch covering algorithm using affine projection method. Mathematics in Computer Science, 12 (2):235–246, June 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Martinez-Moro:2018:F

[402] Edgar Martínez-Moro, Josep M. Miret, and Luis Ramiro Piñeiro. Foreword. Mathematics in Computer Science, 12 (3):247–250, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-018-0371-3.pdf.

Badia:2018:EPM

[403] Valentina Badia, Hebert Pérez-Rosés, and Joe Ryan. Eulogy for Professor Mirka Miller (1949–2016). Mathematics in Computer Science, 12(3):251– 254, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Ryan:2018:BSC

[404] Joe Ryan. A brief survey on the contribution of Mirka Miller to the security of statistical databases. Mathematics in Computer Science, 12(3):255– 262, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Salas:2018:SBP

[405] Julián Salas and Josep Domingo-Ferrer. Some basics on privacy techniques, anonymization and their big data challenges. Mathematics in Computer Science, 12(3):263–274, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Perez-Roses:2018:SYN

[406] Hebert Pérez-Rosés. Sixty years of network reliability. Mathematics in Computer Science, 12(3):275–293, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Abarzua:2018:ASC

[407] Rodrigo Abarzúa, Santi Martínez, Valeria Mendoza, and Javier Valera. Avoiding side-channel attacks by computing isogenous and isomorphic elliptic curves. Mathematics in Computer Science, 12 (3):295–307, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Miret:2018:PBC

[408] Josep M. Miret, Daniel Sadornil, and Juan G. Tena. Pairing-based cryptography on elliptic curves. Mathematics in Computer Science, 12(3):309– 318, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Adj:2018:STP

[409] Gora Adj, Isaac Canales-Martínez, Luis Rivera-Zamarripa, and Francisco Rodríguez-Henríquez. Smoothness test for polynomials defined over small characteristic finite fields. Mathematics in Computer Science, 12(3):319– 337, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Borges-Quintana:2018:WOI

[410] Mijail Borges-Quintana, Miguel Ángel Borges-Trenard, and Edgar Martínez-Moro. On the weak order ideal associated to linear codes. Mathematics in Computer Science, 12(3):339– 347, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Climent:2018:BFD

[411] Joan-Josep Climent, Francisco J. García, and Verónica Requena. Boolean functions: Degree and support. Mathematics in Computer Science, 12(3):349– 369, September 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Goyeneche:2018:F

[412] Dardo Goyeneche, Ilias Kotsireas, and Padraig Ó Catháin. Foreword. Mathematics in Computer Science, 12(4): 371–372, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-018-0388-7.pdf.

Djokovic:2018:GSD

[413] Dragomir Z. – Doković and Ilias S. Kotsireas. Goethals–Seidel difference families with symmetric or skew base blocks. Mathematics in Computer Science, 12 (4):373–388, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Crnkovic:2018:NST

[414] Dean Crnković and Ronan Egan. A note on Siamese twin designs intersecting in a BIBD and a PBD. Mathematics in Computer Science, 12(4):389–395, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Acevedo:2018:RDS

[415] Santiago Barrera Acevedo and Heiko Dietrich. Relative difference sets and Hadamard matrices from perfect quaternionic arrays. Mathematics in Computer Science, 12(4):397–406, December 2018.

CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Alvarez:2018:MHG

[416] V. Alvarez, J. A. Armario, R. M. Falcón, M. D. Frau, F. Gudiel, M. B. Güemes, and A. Osuna. A mixed heuristic for generating cocyclic Hadamard matrices. Mathematics in Computer Science, 12 (4):407–417, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Armario:2018:QHF

[417] José Andrés Armario, Iván Bailera, Joaquim Borges, and Josep Rifà. Quasi-Hadamard full propelinear codes. Mathematics in Computer Science, 12(4): 419–428, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Colbourn:2018:CFF

[418] Charles J. Colbourn and Violet R. Syrotiuk. On a combinatorial framework for fault characterization. Mathematics in Computer Science, 12(4):429– 451, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Gillespie:2018:COA

[419] Neil I. Gillespie, Padraig Ó Catháin, and Cheryl E. Praeger. Construction of the outer automorphism of S_{6} via a complex Hadamard matrix. Mathematics in Computer Science, 12(4): 453–458, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-018-0382-0.pdf.

Bruzda:2018:ESC

[420] Wojciech T. Bruzda. Extension of the set of complex Hadamard matrices of size 8. Mathematics in Computer Science, 12(4):459–464, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-018-0379-8.pdf.

Winterhof:2018:NES

[421] Arne Winterhof, Oguz Yayla, and Volker Ziegler. Non-existence of some nearly perfect sequences, near Butson– Hadamard matrices, and near conference matrices. Mathematics in Computer Science, 12(4):465–471, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Rajchel:2018:RHM

[422] Grzegorz Rajchel, Adam Gasiorowski, and Karol Zyczkowski. Robust Hadamard matrices, unistochastic rays in Birkhoff polytope and equi-entangled bases in composite spaces. Mathematics in Computer Science, 12(4): 473–490, December 2018. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-018-0384-y.pdf.

Beaudin:2019:F

[423] Michel Beaudin, Thierry Dana-Picard, Alexander Levin, Christoph Koutschan, Ilias Kotsireas, and Daniel Robertz. Foreword. Mathematics in Computer Science, 13(1–2):1–3, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-019-00391-x.pdf.

Benjamin:2019:APS

[424] Jurell Benjamin, Donna Walker, Aleksandr Mylläri, and Tatiana Mylläri. On the applicability of pairwise separations method in astronomy: Influence of the noise in data. Mathematics in Computer Science, 13(1–2):5–10, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Blanco-Trejo:2019:FTA

[425] S. Blanco-Trejo, C. Alemán-Morillo, F. Díaz del Río, and P. Real. Fractal topological analysis for 2D binary digital images. Mathematics in Computer Science, 13(1–2):11–20, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Blaser:2019:DC

[426] Nello Blaser and Morten Brun. Divisive cover. Mathematics in Computer Science, 13(1–2):21–29, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-018-0352-6.pdf.

Blazek:2019:LCD

[427] Jirí Blazek and Pavel Pech. Locus computation in dynamic geometry environment. Mathematics in Computer Science, 13(1–2):31–40, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Chan:2019:MHC

[428] E. Y. S. Chan and R. M. Corless. Minimal height companion matrices for Euclid polynomials. Mathematics in Computer Science, 13(1–2):41–56, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Dana-Picard:2019:ASR

[429] Thierry Dana-Picard. Automated study of a regular trifolium. Mathematics in Computer Science, 13(1–2):57–67, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Durcheva:2019:HUC

[430] Mariana Durcheva. How to use CAS (Maple) to help students learn number theory. Mathematics in Computer Science, 13(1–2):69–78, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Fukasaku:2019:MHQ

[431] Ryoya Fukasaku, Hidenao Iwane, and Yosuke Sato. On multivariate Hermitian quadratic forms. Mathematics in Computer Science, 13(1–2):79–93, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Hasek:2019:DGS

[432] R. Hasek. Dynamic geometry software supplemented with a computer algebra system as a proving tool. Mathematics in Computer Science, 13(1–2):95–104, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kagan:2019:PPR

[433] Mikhail Kagan and Brian Mata. A physics perspective on the resistance distance for graphs. Mathematics in Computer Science, 13(1–2):105–115, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Karsenty:2019:CAC

[434] Avi Karsenty and Yaakov Mandelbaum. Computer algebra challenges in nanotechnology: Accurate modeling of nanoscale Electro-optic devices using finite elements method. Mathematics in Computer Science, 13(1–2):117–130, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kovacs:2019:ACA

[435] Zoltán Kovács. Achievements and challenges in automatic locus and envelope animations in dynamic geometry. Mathematics in Computer Science, 13 (1–2):131–141, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-018-0390-0.pdf.

Kozera:2019:MHI

[436] R. Kozera and M. Wilkołazka. A modified Hermite interpolation with exponential parameterization. Mathematics in Computer Science, 13(1– 2):143–155, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-018-0362-4.pdf.

Levin:2019:BDQ

[437] Alexander Levin. Bivariate dimension quasi-polynomials of difference– differential field extensions with weighted basic operators. Mathematics in Computer Science, 13(1–2):157–168, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Mansour:2019:PPC

[438] Toufik Mansour and Matthias Schork.

Permutation patterns and cell decompositions. Mathematics in Computer Science, 13(1–2):169–183, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Nabeshima:2019:SPI

[439] Katsusuke Nabeshima and Shinichi Tajima. Solving parametric ideal membership problems and computing integral numbers in a ring of convergent power series via comprehensive Gröbner systems. Mathematics in Computer Science, 13(1–2):185–194, June 2019. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Naiman:2019:AFA

[440] Aaron E. Naiman. Automated function analysis for calculus. Mathematics in Computer Science, 13(1–2):195–204, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Ohara:2019:ACG

[441] Katsuyoshi Ohara and Shinichi Tajima. An algorithm for computing Grothendieck local residues I: Shape basis case. Mathematics in Computer Science, 13(1–2): 205–216, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Ovodenko:2019:UDT

[442] Regina Ovodenko and Anatoli Kouropatov. The use of digital tools to confront errors during advanced calculus learning: The case of the inflection point. Mathematics in Computer Science, 13 (1–2):217–236, June 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Galan-Garcia:2019:F

[443] José L. Galán-García, Gabriel Aguilera-Venegas, and María Á. Galán-García.

Foreword. Mathematics in Computer Science, 13(3):329–331, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-019-00408-5.pdf.

Lebedev:2019:MCQ

[444] Eugene Lebedev and Hanna Livinska.

Multi-channel queueing networks with input flow controlled by semi–Markov process. Mathematics in Computer Science, 13(3):333–340, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kohlwey:2019:HEE

[445] Elena Kohlwey and Melven Röhrig-Zöllner. Half-explicit exponential Runge–Kutta methods for index-1 DAEs in helicopter simulation. Mathematics in Computer Science, 13(3):341– 365, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Siggel:2019:TOS

[446] Martin Siggel, Jan Kleinert, Tobias Stollenwerk, and Reinhold Maierl. TiGL: An open source computational geometry library for parametric aircraft design. Mathematics in Computer Science, 13 (3):367–389, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Juarez-del-Toro:2019:AIC

[447] R. Juarez del Toro, J. G. Castrejón-Lozano, C. A. Gomez-Rosales, and S. López-Chavarría. Application of an intelligent control on economics dynamic system: The attractive invariant ellipsoid approach. Mathematics in Computer Science, 13(3):391–401, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

deLeon:2019:CBM

[448] José Alfredo Sánchez de León. Calculation of binomial and multinomial coefficients by sequences of summations. Mathematics in Computer Science, 13 (3):403–415, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Roanes-Lozano:2019:SRA

[449] Eugenio Roanes-Lozano, Jose Luis Galán-García, and Carmen Solano-Macías. Some reflections about the success and impact of the computer algebra system DERIVE with a 10-year time perspective. Mathematics in Computer Science, 13(3):417–431, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Aydogan:2019:CFC

[450] S. Melike Aydogan and F. Müge Sakar. On convex functions with complex order through bounded boundary rotation. Mathematics in Computer Science, 13 (3):433–439, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Sakar:2019:BIC

[451] F. Müge Sakar and S. Melike Aydogan. Bounds on initial coefficients for a certain new subclass of bi-univalent functions by means of Faber polynomial expansions. Mathematics in Computer Science, 13(3):441–447, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Roanes-Lozano:2019:AAD

[452] Eugenio Roanes-Lozano, Rubén González-Martín, and Javier Montero. An algebraic approach to DC railway electrification verification. Mathematics in Computer Science, 13(3):449–457, September 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Davenport:2019:F

[453] James H. Davenport, Laura Kovacs, and Daniela Zaharie. Foreword. Mathematics in Computer Science, 13(4): 459–460, December 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-019-00411-w.pdf.

Huang:2019:UML

[454] Zongyan Huang, Matthew England, David J. Wilson, James Bridge, James H. Davenport, and Lawrence C. Paulson. Using machine learning to improve cylindrical algebraic decomposition. Mathematics in Computer Science, 13(4):461–488, December 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-019-00394-8.pdf.

Andrei:2019:PBE

[455] Stefan Andrei, Albert M. K. Cheng, and Vlad Radulescu. Processor bounding for an efficient non-preemptive task scheduling algorithm. Mathematics in Computer Science, 13(4):489–515, December 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Lichtblau:2019:APG

[456] Daniel Lichtblau. Approximate polynomial GCD by approximate syzygies. Mathematics in Computer Science, 13 (4):517–532, December 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Heule:2019:OSB

[457] Marijn J. H. Heule. Optimal symmetry breaking for graph problems. Mathematics in Computer Science, 13(4):533– 548, December 2019. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Panayotopoulos:2020:MC

[458] A. Panayotopoulos. On meandric colliers. Mathematics in Computer Science, 14(1):1–8, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

daSilva:2020:VCG

[459] Eduardo Sant’Ana da Silva and Helio Pedrini. Vertex coloring of a graph for memory constrained scenarios. Mathematics in Computer Science, 14(1):9–17, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Decker:2020:MTN

[460] Wolfram Decker, Christian Eder, Viktor Levandovskyy, and Sharwan K. Tiwari. Modular techniques for noncommutative Gröbner bases. Mathematics in Computer Science, 14(1):19–33, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Sari:2020:DCS

[461] Mustafa Sari and Emre Kolotoglu. A different construction for some classes of quantum MDS codes. Mathematics in Computer Science, 14(1):35–44, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Dana-Picard:2020:EIF

[462] Thierry Dana-Picard, Aharon Naiman, Witold Mozgawa, and Waldemar Cieślak. Exploring the isoptics of Fermat curves in the affine plane using DGS and CAS. Mathematics in Computer Science, 14(1):45–67, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Blecher:2020:DC

[463] Aubrey Blecher, Charlotte Brennan, Arnold Knopfmacher, and Toufik Mansour. The depth of compositions. Mathematics in Computer Science, 14(1): 69–76, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Manin:2020:NDP

[464] Yuri I. Manin and Matilde Marcolli.

Nori diagrams and persistent homology. Mathematics in Computer Science, 14 (1):77–102, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Aksim:2020:EAB

[465] Dan Aksim and Dmitry Pavlov. On the extension of Adams–Bashforth–Moulton methods for numerical integration of delay differential equations and application to the Moon’s orbit. Mathematics in Computer Science, 14(1):103–109, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Dey:2020:THR

[466] Papri Dey and Daniel Plaumann. Testing hyperbolicity of real polynomials. Mathematics in Computer Science, 14 (1):111–121, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Hashemi:2020:CCB

[467] Amir Hashemi, Martin Kreuzer, and Samira Pourkhajouei. Computing coupled border bases. Mathematics in Computer Science, 14(1):123–140, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Choudhary:2020:CTG

[468] Aruni Choudhary, Siargey Kachanovich, and Mathijs Wintraecken. Coxeter triangulations have good quality. Mathematics in Computer Science, 14 (1):141–176, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-020-00461-5.pdf.

Horacek:2020:SBB

[469] Jan Horácek, Martin Kreuzer, and Ange-Salomé Messeng Ekossono. A signature based border basis algorithm. Mathematics in Computer Science, 14 (1):177–189, March 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Bigatti:2020:F

[470] Anna Maria Bigatti, Francisco Botana, Thierry Dana-Picard, Felipe Gago, Ilias Kotsireas, Manuel Ladra, and Wei Li. Foreword. Mathematics in Computer Science, 14(2):191–192, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-019-00445-0.pdf.

Almech:2020:AGD

[471] Alberto Almech and Eugenio Roanes-Lozano. Automatic generation of diagrammatic subway maps for any date with Maple. Mathematics in Computer Science, 14(2):193–207, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Ku-Cauich:2020:CER

[472] Juan Carlos Ku-Cauich and Guillermo Morales-Luna. Conversion of element representations in Galois rings. Mathematics in Computer Science, 14(2): 209–222, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Kozera:2020:NMH

[473] R. Kozera and M. Wilkołazka. A note on modified Hermite interpolation. Mathematics in Computer Science, 14(2): 223–239, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Minglibayev:2020:CPT

[474] Mukhtar Minglibayev, Alexander Prokopenya, and Saule Shomshekova. Computing perturbations in the two-planetary three-body problem with masses varying non-isotropically at different rates. Mathematics in Computer Science, 14 (2):241–251, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Wojas:2020:FSD

[475] Włodzimierz Wojas, Jan Krupa, and Jarosław Bojarski. Familiarizing students with definition of Lebesgue outer measure using Mathematica: Some examples of calculation directly from its definition. Mathematics in Computer Science, 14(2):253–270, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-019-00435-2.pdf.

Prank:2020:ACE

[476] Rein Prank. Analysing the “calculator effect” of different kinds of software for school arithmetic and algebra. Mathematics in Computer Science, 14 (2):271–279, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Braun:2020:NAV

[477] Elishan Braun, Werner M. Seiler, and Matthias Seiß. On the numerical analysis and visualisation of implicit ordinary differential equations. Mathematics in Computer Science, 14(2):281–293, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Edneral:2020:NOS

[478] Victor F. Edneral and Alexander G. Petrov. Nonlinear oscillations of a spring pendulum at the 1:1:2 resonance by normal form methods. Mathematics in Computer Science, 14(2):295–303, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Perminov:2020:IHD

[479] A. S. Perminov and E. D. Kuznetsov. The implementation of Hori–Deprit method to the construction averaged planetary motion theory by means of computer algebra system Piranha. Mathematics in Computer Science, 14 (2):305–316, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Bavula:2020:CSM

[480] V. V. Bavula. Classification of simple modules of the Ore extension K[X][Y; f \frac{d}{dX}]. Mathematics in Computer Science, 14(2):317–325, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-019-00414-7.pdf.

Falkensteiner:2020:SFO

[481] Sebastian Falkensteiner and J. Rafael Sendra. Solving first order autonomous algebraic ordinary differential equations by places. Mathematics in Computer Science, 14(2):327–337, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Bavula:2020:SCA

[482] V. V. Bavula. Skew category algebras. Mathematics in Computer Science, 14(2):339–346, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-019-00415-6.pdf.

Evgrafov:2020:DPE

[483] Alexander Evgrafov and Alexander Levin. Dimension polynomials and the Einstein’s strength of some systems of quasi-linear algebraic difference equations. Mathematics in Computer Science, 14(2):347–360, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Levin:2020:MDD

[484] Alexander Levin. Multivariate difference– differential dimension polynomials. Mathematics in Computer Science, 14 (2):361–374, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Falcon:2020:UCD

[485] Raúl M. Falcón. Using a CAS/DGS to analyze computationally the configuration of planar bar linkage mechanisms based on partial latin squares. Mathematics in Computer Science, 14 (2):375–389, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Hasek:2020:EDC

[486] Roman Hasek. Exploration of dual curves using a dynamic geometry and computer algebra system. Mathematics in Computer Science, 14(2):391–398, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Takato:2020:UOS

[487] Setsuo Takato and José A. Vallejo. Using Oshima splines to produce accurate numerical results and high quality graphical output. Mathematics in Computer Science, 14(2):399–413, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Bernal:2020:CAC

[488] José Joaquín Bernal, Diana H. Bueno-Carreño, and Juan Jacobo Simón. Constructions of Abelian codes multiplying dimension of cyclic codes. Mathematics in Computer Science, 14(2):415–421, June 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Bouyuklieva:2020:BIC

[489] Stefka Bouyuklieva, Radka Russeva, and Emine Karatash. Binary isodual codes having an automorphism of odd prime order. Mathematics in Computer Science, 14(2):423–429, June 2020. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Labahn:2020:F

[490] George Labahn, James H. Davenport, and Josef Urban. Foreword. Mathematics in Computer Science, 14(3): 531–532, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-020-00475-z.pdf.

Ekici:2020:MLC

[491] Burak Ekici and Cezary Kaliszyk. Mac Lane’s comparison theorem for the Kleisli construction formalized in Coq. Mathematics in Computer Science, 14 (3):533–549, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-020-00450-8.pdf.

Lewis:2020:IAI

[492] Robert H. Lewis. Image analysis: Identification of objects via polynomial systems. Mathematics in Computer Science, 14(3):551–558, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-020-00451-7.pdf.

Bercic:2020:DFD

[493] Katja Bercic and Janos Vidali. Discrete-ZOO: A fingerprint database of discrete objects. Mathematics in Computer Science, 14(3):559–575, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-020-00453-5.pdf.

Brysiewicz:2020:NSC

[494] Taylor Brysiewicz. Numerical software to compute Newton polytopes and tropical membership. Mathematics in Computer Science, 14(3): 577–589, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-020-00454-4.pdf.

Alpuente:2020:ESE

[495] M. Alpuente, D. Ballis, and J. Sapiña. Efficient safety enforcement for Maude programs via program specialization in the ÁTAME system. Mathematics in Computer Science, 14(3): 591–606, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-020-00455-3.pdf.

Kim:2020:PSP

[496] Seyeon Kim, Marco Pollanen, Michael G. Reynolds, and Wesley S. Burr. Problem solving as a path to comprehension. Mathematics in Computer Science, 14 (3):607–621, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-020-00457-1.pdf.

Jiu:2020:CPZ

[497] Lin Jiu and Christoph Koutschan. Calculation and properties of zonal polynomials. Mathematics in Computer Science, 14(3):623–640, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-020-00458-0.pdf.

DiCrescenzo:2020:ESD

[498] Giovanni Di Crescenzo, Matluba Khodjaeva, Delaram Kahrobaei, and Vladimir Shpilrain. Efficient and secure delegation of exponentiation in general groups to a single malicious server. Mathematics in Computer Science, 14(3): 641–656, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-020-00462-4.pdf.

Jakubuv:2020:RWP

[499] Jan Jakubuv and Cezary Kaliszyk. Relaxed weighted path order in theorem proving. Mathematics in Computer Science, 14(3):657–670, September 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

link.springer.com/content/pdf/10.
1007/s11786-020-00474-0.pdf.

Chen:2020:FSF

[500] Xiaoyu Chen, Hongbo Li, and Jing Yang. Foreword to the special focus on automated deduction in geometry. Mathematics in Computer Science, 14 (4):671–672, December 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-020-00488-8.pdf.

Quaresma:2020:ADK

[501] Pedro Quaresma. Automated deduction and knowledge management in geometry. Mathematics in Computer Science, 14(4):673–692, December 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

                                        springer.com/content/pdf/10.1007/
                                        s11786-020-00489-7.pdf.

Todd:2020:SDG

[502] Philip Todd. A symbolic dynamic geometry system using the analytical geometry method. Mathematics in Computer Science, 14(4):693–726, December 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-020-00490-0.pdf.

Kovacs:2020:ADI

[503] Zoltán Kovács. Automated detection of interesting properties in regular polygons. Mathematics in Computer Science, 14(4):727–755, December 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://

                                        link.springer.com/content/pdf/10.
                                        1007/s11786-020-00491-z.pdf.

Wang:2020:SAA

[504] Dongming Wang, Bo Huang, and Xiaoyu Chen. On n-sectors of the angles of an arbitrary triangle. Mathematics in Computer Science, 14(4): 757–773, December 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-020-00492-y.pdf.

Todd:2020:SAD

[505] Philip Todd. A system for automated deduction in engineering mechanics. Mathematics in Computer Science, 14 (4):775–790, December 2020. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-020-00493-x.pdf.

Loja:2021:FSF

[506] Amélia Loja, Paulo Vasconcelos, Joaquim Infante Barbosa, and José Alberto Rodrigues. Foreword to the special focus on advances in symbolic and numeric computation III. Mathematics in Computer Science, 15(1):1–4, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http://link.

springer.com/content/pdf/10.1007/
s11786-020-00487-9.pdf.

daRocha:2021:CPB

[507] Zélia da Rocha. Common points between perturbed Chebyshev polynomials of second kind. Mathematics in Computer Science, 15(1):5–13, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Mesquita:2021:OPS

[508] Teresa Augusta Mesquita. On a 2-orthogonal polynomial sequence via quadratic decomposition. Mathematics in Computer Science, 15(1):15–31, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Escobar:2021:INT

[509] J. M. Escobar, J. Núñez-Valdés, and P. Pérez-Fernández. Introducing a new two-parameter invariant function for algebras. Mathematics in Computer Science, 15(1):33–44, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Matos:2021:EFI

[510] José M. A. Matos, Maria João Rodrigues, and João Carrilho de Matos. Explicit formulae for integro–differential operational matrices. Mathematics in Computer Science, 15(1):45–61, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Conceicao:2021:SCA

[511] Ana C. Conceição. Symbolic computation applied to the study of the kernel of special classes of paired singular integral operators. Mathematics in Computer Science, 15(1):63–90, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Ammi:2021:GSC

[512] Moulay Rchid Sidi Ammi, Mostafa Tahiri, and Delfim F. M. Torres. Global stability of a Caputo fractional SIRS model with general incidence rate. Mathematics in Computer Science, 15 (1):91–105, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Silva:2021:SSO

[513] Cristiana J. Silva and Guillaume Cantin. Synchronization and self-organization in complex networks for a tuberculosis model. Mathematics in Computer Science, 15(1):107–120, March 2021. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Mota:2021:EEB

[514] P. R. Mota and P. B. Vasconcelos. Estimating the employment band of inaction with multiple breaks due to labor market reforms. Mathematics in Computer Science, 15(1):121–133, March 2021. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Adak:2021:CVE

[515] Dibyendu Adak and Sundararajan Natarajan. On the H^{1} conforming virtual element method for time dependent Stokes equation. Mathematics in Computer Science, 15(1):135–154, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Piloto:2021:CST

[516] Paulo A. G. Piloto, Carlos Balsa, Fernando Ribeiro, and Ronaldo Rigobello. Computational simulation of the thermal effects on composite slabs under fire conditions. Mathematics in Computer Science, 15(1):155–171, March 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

England:2021:FDA

[517] Matthew England, Wolfram Koepf, and Thomas Sturm. Foreword, with a dedication to Andreas Weber. Mathematics in Computer Science, 15(2):173–175, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00476-
y.

Boulier:2021:SCT

[518] François Boulier, François Lemaire, and Adrien Poteaux. A short contribution to the theory of regular chains. Mathematics in Computer Science, 15(2):177–188, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00477-
x.

Cano:2021:ARP

[519] José Cano, Sebastian Falkensteiner, and J. Rafael Sendra. Algebraic, rational and Puiseux series solutions of systems of autonomous algebraic ODEs of dimension one. Mathematics in Computer Science, 15(2):189–198, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00478-
w.

Grigoriev:2021:EER

[520] Dima Grigoriev, Alexandru Iosif, and Andreas Weber. Efficiently and effectively recognizing toricity of steady state varieties. Mathematics in Computer Science, 15(2):199–232, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00479-
9.

Hashemi:2021:DUB

[521] Amir Hashemi, Hossein Parnian, and Werner M. Seiler. Degree upper bounds for involutive bases. Mathematics in Computer Science, 15(2):233–254, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00480-
2.

Hu:2021:TDG

[522] Youren Hu and Xiao-Shan Gao. Tropical differential Gröbner bases. Mathematics in Computer Science, 15(2):255–269, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00481-
1.

Imbach:2021:CCZ

[523] Rémi Imbach, Marc Pouget, and Chee Yap. Clustering complex zeros of triangular systems of polynomials. Mathematics in Computer Science, 15(2): 271–292, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-020-
00482-0.

Mohammadi:2021:SBA

[524] Zahra Mohammadi, Gregory J. Reid, and S.-L. Tracy Huang. Symmetrybased algorithms for invertible mappings of polynomially nonlinear PDE to linear PDE. Mathematics in Computer Science, 15(2):293–316, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00483-
z.

Nabeshima:2021:TZD

[525] Katsusuke Nabeshima and Shinichi Tajima. Testing zero-dimensionality of varieties at a point. Mathematics in Computer Science, 15(2):317–331, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00484-
y.

Seiler:2021:LBA

[526] Werner M. Seiler, Matthias Seiß, and Thomas Sturm. A logic based approach to finding real singularities of implicit ordinary differential equations. Mathematics in Computer Science, 15(2):333–352, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00485-
x.

Tajima:2021:ACT

[527] Shinichi Tajima and Katsusuke Nabeshima. An algorithm for computing torsion differential forms associated with an isolated hypersurface singularity. Mathematics in Computer Science, 15(2): 353–367, June 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-020-
00486-w.

England:2021:FDV

[528] Matthew England, François Boulier, and Thomas Sturm. Foreword, with a dedication to Vladimir Gerdt. Mathematics in Computer Science, 15(3):369– 371, September 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-021-
00509-0.

Brehard:2021:SNV

[529] Florent Bréhard. A symbolic-numeric validation algorithm for linear ODEs with Newton–Picard method. Mathematics in Computer Science, 15(3):373– 405, September 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-021-
00510-7.

Gutnik:2021:SCE

[530] Sergey A. Gutnik and Vasily A.

Sarychev. Symbolic computations of the equilibrium orientations of a system of two connected bodies moving on a circular orbit around the Earth. Mathematics in Computer Science, 15(3):407–417, September 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00511-
6.

Hashemi:2021:IGA

[531] Amir Hashemi, Thomas Izgin, and Werner M. Seiler. An involutive GVW algorithm and the computation of Pommaret bases. Mathematics in Computer Science, 15(3):419–452, September 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00512-
5.

Hashemi:2021:RGI

[532] Amir Hashemi, Matthias Orth, and Werner M. Seiler. Relative Gröbner and involutive bases for ideals in quotient rings. Mathematics in Computer Science, 15(3):453–482, September 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11786-021-00513-
                                        4.

Koutschan:2021:CTM

[533] Christoph Koutschan and Elaine Wong. Creative telescoping on multiple sums. Mathematics in Computer Science, 15 (3):483–498, September 2021. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11786-021-00514-3.

Kruff:2021:ARB

[534] Niclas Kruff, Christoph Lüders, and Sebastian Walcher. Algorithmic reduction of biological networks with multiple time scales. Mathematics in Computer Science, 15(3):499–534, September 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-021-00515-
                                        2.

Ferras:2021:HOM

[535] Luís L. Ferrás, Neville Ford, and Magda Rebelo. High-order methods for systems of fractional ordinary differential equations and their application to time-fractional diffusion equations. Mathematics in Computer Science, 15(4):535–551, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-019-00448-
                                        x.

Beltrametti:2021:GHT

[536] M. C. Beltrametti, C. Campi, and M. Torrente. Geometry of the Hough transforms with applications to synthetic data. Mathematics in Computer Science, 15(4):553–575, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00470-
4.

Schneider:2021:AMP

[537] Carsten Schneider. The absentminded passengers problem: a motivating challenge solved by computer algebra. Mathematics in Computer Science, 15(4):577–588, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00494-
w.

Middeke:2021:CFF

[538] Johannes Middeke, David J. Jeffrey, and Christoph Koutschan. Common factors in fraction-free matrix decompositions. Mathematics in Computer Science, 15(4):589–608, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-020-00495-
9.

Berkemer:2021:CPA

[539] Sarah J. Berkemer, Christian Höner zu Siederdissen, and Peter F. Stadler. Compositional properties of alignments. Mathematics in Computer Science, 15 (4):609–630, December 2021. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL https:

//link.springer.com/article/10.
1007/s11786-020-00496-8.

Botana:2021:MG

[540] Francisco Botana, Zoltán Kovács, and Tomás Recio. A mechanical geometer. Mathematics in Computer Science, 15(4):631–641, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-020-00497-
                                        7.

Ortegaray:2021:HKA

[541] Andrew Ortegaray, Robert C. Berwick, and Matilde Marcolli. Heat kernel analysis of syntactic structures. Mathematics in Computer Science, 15(4):643– 660, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11786-021-
                                        00498-0.

Solin:2021:SPI

[542] Pavel Solin. Self-paced, instructorassisted approach to teaching linear algebra. Mathematics in Computer Science, 15(4):661–687, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-021-00499-
                                        z.

Roanes-Lozano:2021:UFE

[543] Eugenio Roanes-Lozano and Carmen Solano-Macías. Using fractals and turtle geometry to visually explain the spread of a virus to kids: a STEM multitarget activity. Mathematics in Computer Science, 15(4):689–699, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11786-021-00500-
9.

Mansour:2021:CLP

[544] Toufik Mansour, José L. Ramírez, and Diana A. Toquica. Counting lattice points on bargraphs of Catalan words. Mathematics in Computer Science, 15(4):701–713, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00501-
8.

Kacker:2021:FEC

[545] Raghu N. Kacker, D. Richard Kuhn, and Dimitris E. Simos. Factorials experiments, covering arrays, and combinatorial testing. Mathematics in Computer Science, 15(4):715–739, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00502-

7. See correction [546].

Kacker:2021:CFE

[546] Raghu N. Kacker, D. Richard Kuhn, and Dimitris E. Simos. Correction to: Factorials experiments, covering arrays, and combinatorial testing. Mathematics in Computer Science, 15(4):741, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00516-

1. See [545].

Jia:2021:LEK

[547] Nan Jia, Yaping Mao, and Eddie Cheng. Linear k-arboricity of caylay graphs on Abelian groups with given degree. Mathematics in Computer Science, 15(4):743–755, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00503-
6.

Dana-Picard:2021:EOT

[548] Thierry Dana-Picard. Envelopes and offsets of two algebraic plane curves: Exploration of their similarities and differences. Mathematics in Computer Science, 15(4):757–774, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00504-
5.

Aggarwal:2021:XBF

[549] Anurag Aggarwal, Zoltán Kovács, and Jonathan Wolfe. XaoS 4.0 and beyond: Fractals in mathematics education. Mathematics in Computer Science, 15(4):775–788, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00505-
4.

Kovacs:2021:TAC

[550] Zoltán Kovács. Two almost-circles, and two real ones. Mathematics in Computer Science, 15(4):789–801, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00506-
3.

Shu:2021:PIE

[551] Kevin Shu, Andrew Ortegaray, and Matilde Marcolli. Phylogenetics of Indo-European language families via an algebro-geometric analysis of their syntactic structures. Mathematics in Computer Science, 15(4):803–857, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-021-00507-
2.

Adamou:2021:CTV

[552] Ibrahim Adamou and Bernard Mourrain. Computing the topology of Voronoï diagrams of parallel half-lines. Mathematics in Computer Science, 15(4):859– 876, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-021-
00508-1.

Vlachakis:2021:MMM

[553] Dimitrios Vlachakis and Panayiotis Vlamos. Mathematical multidimensional modelling and structural artificial intelligence pipelines provide insights for the designing of highly specific AntiSARS-CoV2 agents. Mathematics in Computer Science, 15(4):877– 888, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-021-
00517-0.

Mansour:2021:CCC

[554] Toufik Mansour, Reza Rastegar, and Armend Sh. Shabani. On column-convex and convex Carlitz polyominoes. Mathematics in Computer Science, 15(4):

889–898, December 2021. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11786-021-
                                        00518-z.

Archibald:2022:PCC

[555] Margaret Archibald, Aubrey Blecher, and Arnold Knopfmacher. Protected cells in compositions. Mathematics in Computer Science, 16(1):??, March 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-021-00519-
                                        y.

Port:2022:TAS

[556] Alexander Port, Taelin Karidi, and Matilde Marcolli. Topological analysis of syntactic structures. Mathematics in Computer Science, 16(1):??, March 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-021-00520-
                                        5.

Spahn:2022:ESG

[557] George Spahn. Enumerating solutions to grid-based puzzles with a fixed number of rows. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00530-
                                        x.

Laurian:2022:FBW

[558] Azebaze Guimagang Laurian, Fouotsa Emmanuel, El Mrabet Nadia, and Pecha Njiahouo Aminatou. Faster beta Weil pairing on BLS pairing friendly curves with odd embedding degree. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00531-
w.

Ishihara:2022:ELP

[559] Yuki Ishihara. Efficient localization at a prime ideal without producing unnecessary primary components. Mathematics in Computer Science, 16(2–3): ??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-022-
00537-4.

Gutnik:2022:CAM

[560] Sergey A. Gutnik and Vasily A.

Sarychev. Computer algebra methods for searching the stationary motions of the connected bodies system moving in gravitational field. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00535-
6.

England:2022:F

[561] Matthew England, François Boulier, Timur Sadykov, and Thomas Sturm. Foreword. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00533-

Hofstadler:2022:CEC

[562] Clemens Hofstadler, Clemens G. Raab, and Georg Regensburger. Computing elements of certain form in ideals to prove properties of operators. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00536-
                                        5.

Chen:2022:FMP

[563] Tian Chen and Michael Monagan. Factoring multivariate polynomials represented by black boxes: a Maple + C implementation. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00534-
                                        7.

Karpov:2022:DLA

[564] Peter Karpov. Design of low-artifact interpolation kernels by means of computer algebra. Mathematics in Computer Science, 16(2–3):??, September 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00538-
                                        3.

Levin:2022:NTD

[565] Alexander Levin. A new type of difference dimension polynomials. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11786-022-00540-
9.

Greuel:2022:USS

[566] Gert-Martin Greuel, Gerhard Pfister, and Hans Schönemann. Using semicontinuity for standard bases computations. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00539-
2.

Wester:2022:MVG

[567] Michael J. Wester. Memories of Vladimir Gerdt. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00541-
8.

Pommaret:2022:HMS

[568] J.-F. Pommaret. How many structure constants do exist in Riemannian geometry? Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00546-
3.

Kotsireas:2022:PRP

[569] Ilias S. Kotsireas. Personal recollections of Prof. Vladimir P. Gerdt. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11786-022-00545-
                                        4.

Wang:2022:SNR

[570] Dongming Wang. SMS Nanning and RWTH Aachen: In memory of Vladimir Gerdt. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00543-
                                        6.

Seiler:2022:YVG

[571] Werner M. Seiler. 30 years with Vladimir Gerdt. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00544-
                                        5.

Buchberger:2022:VGC

[572] Bruno Buchberger. Vladimir Gerdt: a computer algebra enthusiast. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00542-
                                        7.

Robertz:2022:F

[573] Daniel Robertz and Werner M. Seiler. Foreword. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00550-
                                        7.

Ceria:2022:NGJ

[574] Michela Ceria and Ferdinando Mora.

De Nugis Groebnerialium 7: Janet, Gerdt, Tamari. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00549-
0.

Ceria:2022:ABC

[575] Michela Ceria. Applications of bar code to involutive divisions and a “Greedy” algorithm for complete sets. Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00548-
1.

Brown:2022:CAR

[576] Christopher W. Brown, Zoltán Kovács, Tomás Recio, Róbert Vajda, and M. Pilar Vélez. Is computer algebra ready for conjecturing and proving geometric inequalities in the classroom? Mathematics in Computer Science, 16(4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00532-
9.

Beuchler:2022:RQH

[577] Sven Beuchler, Tim Haubold, and Veronika Pillwein. Recurrences for quadrilateral high-order finite elements. Mathematics in Computer Science, 16 (4):??, December 2022. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.

                                        com/article/10.1007/s11786-022-
                                        00547-2.

Roanes-Lozano:2023:CBM

[578] Eugenio Roanes-Lozano. Can I bring my calculator to the exam? Some reflections on the abstraction level of computer algebra systems. Mathematics in Computer Science, 17(1):??, March 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-022-00551-
                                        6.

Crnkovic:2023:CTE

[579] Dean Crnković, Vedrana Mikulić Crnković, and Andrea Svob. Construction of transitive q-analogs of designs. Mathematics in Computer Science, 17 (1):??, March 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11786-022-
                                        00554-3.

Dana-Picard:2023:AEE

[580] Thierry Dana-Picard and Zoltán Kovács. Automated exploration of envelopes and offsets with networking of technologies. Mathematics in Computer Science, 17 (1):??, March 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11786-022-
                                        00555-2.

Chien:2023:TTD

[581] Bui Van Chien, Gérard H. E. Duchamp, Ngo Quoc Hoan, Hoang Ngoc Minh Vincel, and Nguyen Dinh Vu. Towards a theory of domains for harmonic functions and its symbolic counterpart. Mathematics in Computer Science, 17 (1):??, March 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-022-
00552-5.

Stoichev:2023:UPP

[582] Stoicho D. Stoichev and Mustafa Gezek.

Unitals in projective planes of order 25. Mathematics in Computer Science, 17(1):??, March 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-023-
00556-9.

Tajima:2023:CHM

[583] Shinichi Tajima, Katsusuke Nabeshima, Katsuyoshi Ohara, and Yoko Umeta. Computing holonomic D-modules associated to a family of non-isolated hypersurface singularities via comprehensive Gröbner systems of PBW algebra. Mathematics in Computer Science, 17(1):??, March 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-022-00553-
4.

Bujnowski:2023:ETP

[584] Sławomir Bujnowski, Beata Marciniak, Olutayo Oyeyemi Oyerinde, Zbigniew Lutowski, Adam Flizikowski, and Sebastián García Galan. Equalising the transmission properties of graphmodelled networks by introducing the control of the resources used to transmit information. Mathematics in Computer Science, 17(2):??, June 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11786-023-00559-
                                        6.

Roanes-Lozano:2023:SRA

[585] Eugenio Roanes-Lozano and Carmen Solano-Macías. Some reflections about the success and bibliographic impact of the dynamic geometry system GeoGebra. Mathematics in Computer Science, 17(2):??, June 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-023-00564-
                                        9.

Flizikowski:2023:SAN

[586] Adam Flizikowski, Tomasz Marciniak, Tadeusz A. Wysocki, and Olutayo Oyerinde. Selected aspects of non orthogonal multiple access for future wireless communications. Mathematics in Computer Science, 17(2):??, June 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-023-00561-
                                        y.

Srivastava:2023:HLB

[587] Snehil Srivastava and Panchatcharam Mariappan. Hyperbolic lattice Boltzmann method and discrete Boltzmann method for solid–liquid phase change problem. Mathematics in Computer Science, 17(2):??, June 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-023-00563-
                                        w.

Laszlo:2023:CSB

[588] Michael Laszlo and Sumitra Mukherjee. Counting star–battle configurations. Mathematics in Computer Science, 17(2):??, June 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-023-
00558-7.

Santoprete:2023:SPC

[589] Manuele Santoprete. Some polynomial conditions for cyclic quadrilaterals, tilted kites and other quadrilaterals. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00574-
7.

Kudo:2023:CES

[590] Momonari Kudo, Tasuku Nakagawa, and Tsuyoshi Takagi. Correction to: Efficient search for superspecial hyperelliptic curves of genus four with automorphism group containing C_{6}. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00573-

8. See [592].

England:2023:F

[591] Matthew England, François Boulier, Timur Sadykov, and Thomas Sturm. Foreword. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00565-
8.

Kudo:2023:ESS

[592] Momonari Kudo, Tasuku Nakagawa, and Tsuyoshi Takagi. Efficient search for superspecial hyperelliptic curves of genus four with automorphism group containing C_{6}. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00571-

w. See correction [590].

Baldemir:2023:CMD

[593] Fadime Baldemir and Mesut Sahin. Calculating the minimum distance of a toric code via algebraic algorithms. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00566-
7.

Edneral:2023:ICP

[594] Victor F. Edneral. Integrable cases of the polynomial Liénard-type equation with resonance in the linear part. Mathematics in Computer Science, 17 (3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-023-
00567-6.

Gusev:2023:HIP

[595] Alexander A. Gusev, Galmandakh Chuluunbaatar, Ochbadrakh Chuluunbaatar, Sergue I. Vinitsky, Yuri A.

Blinkov, Algirdas Deveikis, Peter O. Hess, and Luong Le Hai. Hermite interpolation polynomials on parallelepipeds and FEM applications. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00568-
5.

Gutnik:2023:IDT

[596] Sergey A. Gutnik and Vasily A.

Sarychev. Investigation of the dynamics of two connected bodies in the plane of a circular orbit using computer algebra methods. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00569-
4.

Krasikov:2023:SCA

[597] Vitaly A. Krasikov. A survey on computational aspects of polynomial amoebas. Mathematics in Computer Science, 17(3-4):??, December 2023. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-023-00570-
x.

Gakkhar:2024:SSG

[598] Sitanshu Gakkhar and Matilde Marcolli. Syntactic structures and the general Markov models. Mathematics in Computer Science, 18(1):??, March 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11786-023-00575-
                                        6.

Checa:2024:MSS

[599] Carles Checa and Ioannis Z. Emiris. Mixed subdivisions suitable for the greedy Canny–Emiris formula. Mathematics in Computer Science, 18(1):??, March 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-024-00577-
                                        y.

Dana-Picard:2024:ISC

[600] Thierry Dana-Picard. Inflexions of spiric curves: a tale of two tori. Mathematics in Computer Science, 18(1):??, March 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-024-00578-
                                        x.

Baril:2024:LSD

[601] Jean-Luc Baril, Javier F. González, and José L. Ramírez. Last symbol distribution in pattern avoiding Catalan words. Mathematics in Computer Science, 18(1):??, March 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-023-00576-
                                        5.

Lima:2024:LRM

[602] Nilson J. Lima, José A. O. Matos, and Paulo B. Vasconcelos. A low-rank matrix approach to compute polynomial approximations of smooth twodimensional functions. Mathematics in Computer Science, 18(2):??, July 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00581-
2.

Lima:2024:SPD

[603] Nilson J. Lima, José M. A. Matos, and Paulo B. Vasconcelos. Solving partial differential problems with Tau toolbox. Mathematics in Computer Science, 18(2):??, July 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-024-
00580-3.

Mesquita:2024:GFO

[604] Teresa Augusta Mesquita. On a general family of 2-orthogonal polynomial eigenfunctions of a third order differential equation via symbolic computation. Mathematics in Computer Science, 18(2):??, July 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00583-
0.

Mitic:2024:CFE

[605] Peter Mitic. A credibility framework for extreme value-at-risk. Mathematics in Computer Science, 18(2):??, July 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00579-
w.

Henneaux:2024:GMS

[606] David Henneaux, Pierre Schrooyen, Philippe Chatelain, and Thierry Magin. A general methodology for symbolically generating manufactured solutions satisfying prescribed conditions: Application to two-phase flows equations. Mathematics in Computer Science, 18(2):??, July 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00584-
z.

Pakdemirli:2024:SCE

[607] Mehmet Pakdemirli and Ihsan Timucin Dolapci. Solution curves of equations in the differential space. Mathematics in Computer Science, 18(3):??, October 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00592-
z.

delRio:2024:LDP

[608] Tereso del Río and Matthew England. Lessons on datasets and paradigms in machine learning for symbolic computation: a case study on CAD. Mathematics in Computer Science, 18(3):??, October 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00591-
0.

Middeke:2024:ARO

[609] Johannes Middeke, David J. Jeffrey, and Aishat Olagunju. Applications of rational orthogonal matrices. Mathematics in Computer Science, 18(3):??, October 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00590-
1.

Levin:2024:NDG

[610] Alexander Levin. New difference Gröbner bases and bivariate difference dimension polynomials. Mathematics in Computer Science, 18(3):??, October 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00589-
8.

Tajima:2024:DMC

[611] Shinichi Tajima and Katsusuke Nabeshima.

A deterministic method for computing Bertini type invariants of parametric ideals. Mathematics in Computer Science, 18(3):??, October 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-024-00588-
9.

Batkhin:2024:RFS

[612] Alexander Batkhin and Zafar Khaydarov. Resonances and formal stability investigation in Hamiltonian systems. Mathematics in Computer Science, 18 (4):??, December 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11786-024-
00596-9.

Gomes:2024:ADC

[613] Diogo A. Gomes, Friedemann Krannich, Bashayer Majrashi, and Ricardo de Lima Ribeiro. Algorithmic detection of conserved quantities for finitedifference schemes. Mathematics in Computer Science, 18(4):??, December 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-024-00595-
                                        w.

Andraschko:2024:SSU

[614] Bernhard Andraschko, Julian Danner, and Martin Kreuzer. SAT solving using XOR–OR–AND normal forms. Mathematics in Computer Science, 18(4):??, December 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-024-00594-
                                        x.

Arguello:2024:CPC

[615] Anahy Santiago Arguello, Guillaume E. Scholz, and Peter F. Stadler. Core potentials: The consensus segmentation conjecture. Mathematics in Computer Science, 18(4):??, December 2024. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-024-00593-
                                        y.

Li:2025:BDG

[616] Xiaoliang Li and Bo Li. A Bertrand duopoly game with differentiated products reconsidered. Mathematics in Computer Science, 19(1):??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-024-00597-
                                        8.

Ruyong:2025:RFI

[617] Feng Ruyong, Guo Zewang, and Xiong Siting. Rational first integrals of separable differential equations. Mathematics in Computer Science, 19(1):??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-025-00600-
w.

Acevedo:2025:BSC

[618] Santiago Barrera Acevedo, Heiko Dietrich, and Kshitija Vaidya. A block structure for cocyclic Hadamard matrices of order 8p. Mathematics in Computer Science, 19(1):??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-025-00598-
1.

Jimenez-Pastor:2025:CAC

[619] Antonio Jimenez-Pastor, Sonia L.

Rueda, and Rafael Delgado. Computing almost commuting bases of ODOs and Gelfand–Dickey hierarchies. Mathematics in Computer Science, 19(1):??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL https://link.springer.com/
article/10.1007/s11786-025-00601-
9.

Anonymous:2025:JNa

[620] Anonymous. Journal navigation. Mathematics in Computer Science, 19(1): ??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Hasunuma:2025:EG

[621] Toru Hasunuma. Exponentiation of graphs. Mathematics in Computer Science, 19(1):??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-025-00609-
                                        1.

Song:2025:QIL

[622] Jichao Song, Wei Niu, and Le Deng. Qualitative investigation of the Lorenz-84 system using computer algebra methods. Mathematics in Computer Science, 19(1):??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-025-00605-
                                        5.

Faross:2025:GBB

[623] Nicolas Faroß and Simon Schwarz. Gröbner bases for Boolean function minimization. Mathematics in Computer Science, 19(1):??, December 2025. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11786-025-00602-
                                        8.